### v1

# DynVFX: Augmenting Real Videos with Dynamic Content

Abstract.
We present a method for augmenting real-world videos with newly generated dynamic content. Given an input video and a simple user-provided text instruction describing the desired content, our method synthesizes dynamic objects or complex scene effects that naturally interact with the existing scene over time. The position, appearance, and motion of the new content are seamlessly integrated into the original footage while accounting for camera motion, occlusions, and interactions with other dynamic objects in the scene, resulting in a cohesive and realistic output video. We achieve this via a zero-shot, training-free framework that harnesses a pre-trained text-to-video diffusion transformer to synthesize the new content and a pre-trained Vision Language Model to envision the augmented scene in detail. Specifically, we introduce a novel inference-based method that manipulates features within the attention mechanism, enabling accurate localization and seamless integration of the new content while preserving the integrity of the original scene. Our method is fully automated, requiring only a simple user instruction. We demonstrate its effectiveness on a wide range of edits applied to real-world videos, encompassing diverse objects and scenarios involving both camera and object motion [^1] Code will be made publicly available.. Project page: https://dynvfx.github.io/
Text-to-video editing, diffusion models
journal: TOG
Figure 1. DynVFX augments real-world videos with new dynamic content described via simple user-provided text instruction.
1. Introduction
Incorporating computer-generated imagery (CGI) into real-world footage has been a transformative capability in film production, enabling the creation of visual effects that would be difficult or impossible to achieve otherwise. For instance, the seamless integration of CGI characters, such as Gollum in The Lord of the Rings or T-Rex in Jurassic Park, has empowered filmmakers to blend fantastical elements with real-world environments, resulting in immersive storytelling. Inspired by these capabilities, we pose a new creative task: augmenting real-world videos with newly generated dynamic content. Specifically, given an input video and a text prompt describing the desired edit, our goal is to synthesize new dynamic objects or complex scene effects, which naturally interact with the existing scene across time (Fig. 1).
Our task poses several new fundamental challenges. First, the generation must be content-aware, such that the position, appearance, and motion of the synthesized dynamic content integrate naturally with the original scene. This entails synthesizing objects that respect occlusions, maintain appropriate relative size and perspective with respect to the camera position, and realistically interact with other dynamic objects. All of this must be achieved while maintaining the integrity of the original video, ensuring that new content enhances the scene without compromising its authenticity.
Our method leverages a pre-trained text-to-video model without any fine-tuning or additional training. Specifically, given an input video along with a short user instruction describing the new content (e.g., “add a massive whale”), our method produces an edited video where the new content is seamlessly integrated into the original video. We achieve it by estimating the residual to the latent representation of the original video. To envision the edited scene and to identify prominent foreground objects and visual elements, we leverage a vision-language model that translates the user’s instructions into detailed prompts for the text-to-video diffusion model.
As our task requires careful placement of the new content, we propose to steer the localization of the edit through Anchor Extended Attention - incorporate a specific set of keys/values extracted from the original video as additional context to the model. Additionally, to improve the edit harmonization with the original scene, we propose to iteratively update the estimated edit residual.
It is worth noting that our method utilizes a publicly available text-to-video model, which exhibits a significant gap in video generation quality compared to recent state-of-the-art video models. Nevertheless, we observe that within our problem formulation and objective, we can distill from this model surprisingly powerful generative capabilities. We demonstrate the effectiveness of our approach on a variety of edits applied to real-world videos.
To summarize, our work makes the following contributions:
  • We introduce a new task of integrating newly generated dynamic content into real-world videos without relying on the user to provide complex references of the effect (for example, a VFX asset or masks to specify where to locate the VFX).
  • We propose a tuning-free, zero-shot method that enables harmonized content integration while maintaining high fidelity to the original scene
  • We propose an automatic VLM-based evaluation metric tailored for our task, considering multiple factors, including original content preservation, new content harmonization, overall visual quality, and alignment with the edit prompt.
  • We demonstrate state-of-the-art results compared to competing methods, achieving the best trade-off between synthesizing new dynamic elements and maintaining high fidelity to the original content.
2. Related Work
Text-to-Video Models. With the rise of large-scale video-text datasets, there has been significant progress in training text-to-video models using new architectures (Bar-Tal et al., 2024; Polyak et al., 2024). While the foundational architecture of diffusion models has been commonly linked to inflating text-to-image U-Net-based models with temporal layers (Wang et al., 2023a; cerspense, 2023; Guo et al., 2023), recently, a new family of Transformer-based models (OpenAI, 2024; Kong et al., 2025; Yang et al., 2024; Zheng et al., 2024; HaCohen et al., 2024), referred to as Diffusion Transformers (DiTs) (Peebles and Xie, 2023), have gained popularity, as DiTs enhance spatial coherence and enable arbitrary aspect ratio and video-length training. In this work, we utilize a publicly available DiT-based model (Yang et al., 2024), CogVideoX, for augmenting real-world videos with newly generated dynamic content in a zero-shot manner.
Figure 2. Controlling fidelity to the original scene using different extended attention mechanisms. (a-b) SDEdit suffers from the original scene preservation/edit fidelity trade-off. (c-e) Three Extended Attention variants during sampling demonstrate different control levels: Full Extended Attention closely reconstructs the input scene, Masked Extended Attention proves too constrained in overlapping regions despite allowing new content emergence, and our Anchor Extended Attn. achieves optimal results by applying dropout – extending attention only at sparse points within selected regions.
Object Insertion in Images. With the advancement of text-to-image models, techniques for image manipulation leveraging such models have also evolved rapidly. Among these advancements, notable progress has been made in the task of instruction-based image editing. Several works (Zhang et al., 2023c; Sheynin et al., 2023; Brooks et al., 2023; Zhang et al., 2023a) have proposed to directly fine-tune generative models on pairs of original and edited images coupled with user-provided instructions. EmuEdit (Sheynin et al., 2023) leverages a diffusion model trained on a large synthetic dataset to perform various editing tasks guided by task embeddings.
The task of object insertion into images belongs to the same category and can be considered a subfield of instruction-based editing methods. For instance, EraseDraw (Canberk et al., 2024), Paint By Inpaint (Wasserman et al., 2024), and ObjectDrop (Winter et al., 2024) leverage inpainting models to create paired-image datasets, which are then used to fine-tune image editing models. However, extending these approaches to videos presents significant challenges. In particular, generating large-scale instruction-paired video datasets can be prohibitively expensive both in time and computational resources, as it requires substantial manual effort to annotate frames and ensure alignment between textual instructions and video content. This cost and complexity make it challenging to adapt existing image-based methodologies directly to the video domain.
Concurrently to our method, Add-It (Tewel et al., 2024) proposes to manipulate the attention features of a pre-trained text-to-image diffusion model to insert objects into images in a training-free manner. While their method relies on weighted global extended attention, we propose to apply extended attention only to specific regions of the source scene, allowing the generation to focus on essential elements.
Controllable Video Generation. Recently, numerous methods have been developed to incorporate various forms of control signals into video generation pipelines. Several video-to-video methods propose to condition the generation on per-frame spatial maps such as depth maps and edge maps (Wang et al., 2023b; Chen et al., 2023). A line of work (Yatim et al., 2023; Zhao et al., 2023; Park et al., 2024; Jeong et al., 2023; Geyer et al., 2023) proposed to utilize a text-to-video model for the task of motion transfer. Unlike these methods, which are not designed to deviate from the existing structures within a video, our approach focuses on integrating additional dynamic elements into the video.
Recent methods (Bar-Tal et al., 2024; Ma et al., 2024; Zhang et al., 2023b; Zi et al., 2024; Zhou et al., 2023) have explored adapting text-to-video models for video inpainting by conditioning on a masked video and a corresponding binary mask. This setup encourages the model to preserve unmasked information while generating new content in the masked region. All of these methods require user-provided masks — an impractical requirement for integrating complex dynamics as it requires anticipating the placement of dynamic objects (e.g., Fig. 5 jellyfish, tsunami, dinosaurs) ahead of time. In contrast, our method allows for automatic new content localization without any user-provided masks. While a static object can be masked with a simple bounding box, manually defining masks for complex motion or interactions per frame is extremely difficult.
Recently, generative Omnimatte (Lee et al., 2024) proposed a method to automatically decompose a video into object layers and their corresponding effects. However, it is not designed for the task of new content generation, as it allows only for the removal of existing scene elements. VideoDoodles (Yu et al., 2023) combines hand-drawn animations with video footage in a scene-aware manner by tracking a user-provided planar canvas in 3D. However, it does not support the creation of non-planar animations. Our framework includes both localizing dynamic new content and integrating in a scene-aware manner without reliance on user-provided positions.
Language Models for Video Content Creation. Advancements in Vision Language Models (VLMs) have enabled methods to utilize such models in various video-related tasks. Some methods (Chen et al., 2024; Yang et al., 2024) use VLMs to produce detailed video captions from a series of frames, which are then utilized to train text-to-video generative models. Other methods utilize such models for achieving better generation controllability. For instance, VideoDirectorGPT (Lin et al., 2024) utilizes a VLM for multi-scene video generation by training diffusion adapters to incorporate additional conditioning inputs, while LVD (Lian et al., 2023) incorporates layout guidance from the VLM during the sampling process. AutoVFX (Hsu et al., 2024) uses an LLM to generate a video editing program pipeline based on the user instruction. In our work, we employ a VLM as a “VFX assistant” that, based on a short user instruction, provides a comprehensive description of the edited video along with the prominent objects present in the scene.
Professional Software For Video Animation In professional visual effects production, tools such as Autodesk Maya (Autodesk, INC., [n. d.]), Blender (Community, 2018), Unreal Engine (Epic Games, [n. d.]), Adobe After Effects (Christiansen, 2013) and Houdini (SideFX Houdini FX. SideFX, [n. d.]) are widely used for creating and compositing complex visual effects. These tools provide artists with precise control over object modeling, animation, and integration into video footage. While powerful, they require significant expertise, extensive manual intervention, and detailed inputs, such as 3D scene reconstruction or motion tracking. All the aforementioned software count on the user to provide the 3D assets. Even though creating input 3D assets has become an easier task to solve, using new datasets (Deitke et al., 2023; Qiu et al., 2024), or generating assets based on a user prompt (For example, Meshy.ai and Alpha3D), still, 3D physical elements (like fluid or explosion) or global/multi-object effects present a significant challenge. In this work, we take the first steps towards a novice user-friendly workflow.
Figure 3. DynVFX pipeline. Top row: Given an input video $\mathcal{V}_{\text{orig}}$, we apply DDIM inversion (see Sec. 3) and extract spatiotemporal keys and values $[\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}]$ from the original noisy latents. Given the user instruction $\mathcal{P}_{\text{VFX}}$ we instruct the VLM to envision the augmented scene and output the text edit prompt $\mathcal{P}_{\text{comp}}$, prominent object descriptions $\mathcal{O}_{\text{orig}}$ that are used to mask out the extracted keys and values and target object descriptions $\mathcal{O}_{\text{edit}}$. Bottom row: We estimate a residual $\bm{x}_{\text{res}}$ to the original video latent ($\bm{x}_{\text{orig}}$). This is done by iteratively applying SDEdit with our Anchor Extended Attention, segmenting the target objects ($\mathcal{O}_{\text{edit}}$) from the clean result, and updating $\bm{x}_{\text{res}}$ accordingly.
3. Preliminaries
Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are a class of generative models that aim to learn a mapping from noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a data distribution $q$. Starting from a Gaussian i.i.d noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the diffusion model $\Phi$ is applied iteratively through a sequence of denoising steps, ultimately producing a clean output sample $\bm{x}_{0}$.
Recently, a new class of latent text-to-video (T2V) models, built on Diffusion Transformers (DiTs) (Peebles and Xie, 2023), has gained significant popularity. These models comprise multi-modal blocks (MMDiT (Esser et al., 2024)) that allow for joint processing of both text and image modalities, allowing each to inform and refine the other’s representations. To process both modalities together, first, a pre-trained encoder compresses an RGB video $\mathcal{V}$ both spatially and temporally to a latent space. Then, the latent is patchified, and the resulting tokens are concatenated to the text tokens produced by the text encoder (Raffel et al., 2023).
In each MMDiT block, text tokens and spatiotemporal tokens are projected into queries, keys and values using separate sets of weights for each modality, and the sequences of the two modalities are concatenated as a joint input for the attention operation: $\mathbf{Q}=[\mathbf{Q}_{\text{text}},\mathbf{Q}_{\text{spatio}}]$, $\mathbf{K}=[\mathbf{K}_{\text{text}},\mathbf{K}_{\text{spatio}}]$ and $\mathbf{V}=[\mathbf{V}_{\text{text}},\mathbf{V}_{\text{spatio}}]$. The attention (Vaswani et al., 2017) operation computes the affinities between the d-dimensional projections $\mathbf{Q},\mathbf{K}$ to yield the output of the layer:
(1) $\mathbf{A}\cdot\mathbf{V}\ \text{where}\ \mathbf{A}=\texttt{Attention}(\mathbf{Q},\mathbf{K})=\texttt{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)$
To capture inter-token relationships, Rotary Position Embeddings (RoPE) (Su et al., 2021) are applied to the input to the attention operation.
4. Method
Given an input video $\mathcal{V}_{\text{orig}}$ and a textual instruction $\mathcal{P}_{\text{VFX}}$, our goal is to synthesize a new video $\mathcal{V}_{\text{VFX}}$, in which new dynamic elements are seamlessly integrated to the existing scene. We address this task by estimating a residual in the latent space of the text-to-video diffusion model, which is added to the latent of the original video. The final edited video is obtained by decoding their sum with the text-to-video diffusion model’s VAE decoder
Tackling this task requires ensuring that the generated content, such as new objects or effects, adheres to the dynamics of the existing scene. The location and size of the new content must align with the camera motion and the environment, while its actions and movements must appropriately respond to other dynamic objects present in the scene. Our framework (illustrated in Fig. 3) addresses these challenges by incorporating the following key components:
  1. (1)
    VLM as a VFX assistant. We utilize a pre-trained VLM to interpret the user’s instructions, reason about the interactions with the scene’s dynamics, and provide a descriptive prompt for the T2V diffusion model by guiding it to act as ”VFX assistant” via a system prompt containing guidelines in the context of our tasks.
  2. (2)
    Localization via Anchor Extended Attention. To steer the localization of the edit and make it content-aware, we propose to utilize Extended Attention during sampling from the T2V DiT model to a set of keys and values extracted from sparse locations in the original video.
  3. (3)
    Content Harmonization via Iterative refinement. To improve the blending of the generated content with the input video and achieve better harmonization, we iteratively update the estimated residual latent by repeating the sampling process with AnchorExtAttn multiple times, progressively reducing the level of noise added at each step.
4.1. VLM as a VFX assistant
To create a fully automatic framework requiring only a simple user-provided instruction describing the desired content, we incorporate a Vision-Language Model (VLM) into our framework. Specifically, given the user instruction $\mathcal{P}_{\text{VFX}}$, along with evenly spaced keyframes from the original video $\mathcal{V}_{\text{orig}}$, we instruct the VLM (Achiam et al., 2023) to provide a detailed caption - a composition prompt $\mathcal{P}_{\text{comp}}$ describing the new composited scene. While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene. To overcome this, we guide the model to reply in an in-context matter by asking it to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly. In this conversation, VLM will “consult” with a VFX artist about how the new content should be integrated into the scene. Based on their discussion, it provides a caption that describes how the added content fits into the scene. This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment. The guidelines include focusing on (1) spatial and dynamic awareness of existing scene elements, (2) preservation of original scene behaviors, and (3) atmospheric coherence between new and existing content. See SM for more details about how we use the VLM to reason about the new scene integration with in-context reasoning for VFX.
We also use the VLM to obtain a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$. See more details about prompting and mask processing in SM.
4.2. Localization via Anchor Extended Attention
A pivotal aspect of our method is the accurate localization of the new content in the existing scene, ensuring alignment with camera motion, occlusions, and scene depth. While the composition prompt can describe the desired location, naive noising-denoising with this prompt introduces a trade-off: As shown in Fig. 2, using SDEdit (Meng et al., 2022) with a high noising timestep fails to retain the original scene, resulting in misaligned new content. In contrast, a low noising timestep, as illustrated in Fig. 2 (b), limits deviations from the original video.
Figure 4. Ablations. (b) Excluding both AnchorExtAttn and the Iterative refinement process results in significant misalignment with the original scene and poor harmonization (e.g., the size of the puppy relative to the scene and boundary artifacts). (c) Omitting AnchorExtAttn leads to incorrect positioning of the new content. (d) Removing iterative refinement results in poor harmonization. Our full method (e) exhibits good localization and harmonization of the edit
To tackle the localization challenge, we extend the attention module during sampling to include the input video’s corresponding attention features. Specifically, we apply DDIM inversion (Song et al., 2020) to the original video $\mathcal{V}_{\text{orig}}$ and extract the spatio-temporal keys and values $K_{\text{orig}},V_{\text{orig}}$ from the attention module of every block in the network and generation timestep $t$. These keys and values are then used to extend the attention mechanism during sampling with $\mathcal{P}_{comp}$ to control the localization of the edit.
When using all keys and values, the extended attention operation can be expressed as:
(2) $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K_{\text{orig}}],[V_{\text{VFX}},V_{\text{orig}}])$
Interestingly, extending keys and values provides more than just global information, but each feature locally encodes corresponding patches in the video. As shown in Fig. 2 (c), extending the attention to the full set of keys and values is sufficient to achieve an approximate reconstruction of the original video. We hypothesize that this occurs because the same positional embedding is applied to the $K_{\text{orig}}$ as to $K_{\text{VFX}}$, aligning the positional embeddings in spatiotemporal locations. This observation provides strong evidence for the significance of the positional embedding. Due to this, the setup proves to be overly restrictive in adhering to $\mathcal{P}_{\text{comp}}$, as illustrated in Fig. 2(c).
Selective Attention. To overcome this limitation, we propose to restrict the extended attention only to specific positions in the source scene. Specifically, we use a selection function $\mathcal{F}$ to determine which keys and values are retained.
An intuitive strategy is to apply region-based attention with Masked Extended Attention by identifying the most critical regions for preserving scene coherence and extending attention with keys and values within these masked regions $M_{\text{orig}}$, i.e., $\mathcal{F}(\mathcal{A})=M_{\text{orig}}\circ\mathcal{A}$. To get $M_{\text{orig}}$, we ask a VLM to provide a list of foreground objects in the original video $\mathcal{O}_{\text{orig}}$ (typically the most spatially prominent elements within a scene) and then we obtain corresponding masks using a text-based segmentation model. For example, Fig. 2(d), where the masked region includes the horse, illustrates the effect of this on the generated content. As shown, this setup preserves high fidelity to the original scene within the mask $M_{\text{orig}}$ and allows new content to emerge in the unmasked regions.
However, this approach proves too constrained in overlapping regions. To address this limitation, we propose our Anchor Extended Attention:
[…]
:=$\mathcal{F}:=\Biggl{\{}\texttt{Drop}_{FG}(M_{\text{orig}})\cup\texttt{Drop}_{BG}(\sim M_{\text{orig}})~{}\circ~{}\mathcal{A}$
This formulation introduces dropout within the masked regions $M_{\text{orig}}$, generating a sparse set of anchor points that guide the generation while preserving flexibility. This formulation introduces a predominantly content-aware selection of anchors from foreground regions, along with a sparser set of background anchors, to achieve robust and spatially coherent integration of new content (e.g. the horse to enforce the knewly added knight to align with its motion). As demonstrated in Fig. 2(e), this balanced approach offers sufficient flexibility for creative edits while preserving key spatial cues from the original scene.
Figure 5. Sample results of our method. See SM for full vide results.
4.3. Content Harmonization
Our anchor extended attention steers the placement of the new content to align with the original scene. However, it does not guarantee precise pixel-level alignment. As can be seen in Fig. 2(e), the legs of the horse are not perfectly aligned with the original horse. To guarantee pixel-level alignment, a straightforward approach is to extract a mask of the new content $\bm{M}_{VFX}^{rgb}$ from the sampling with AnchorExtAttn (Eq. 3) output $\mathcal{\hat{V}}_{comp}$. More concretely, by applying a text-based segmentation model using the added object description provided by the VLM.
The mask can then be used to replace the pixels outside it with the corresponding pixels from the original video: $\mathcal{V}{\text{comp}}[\sim\bm{M}_{VFX}^{rgb}]=\mathcal{V}{\text{orig}}[\sim\bm{M}_{VFX}^{rgb}]$. While this preserves the unaffected regions, it often results in poor harmonization with the input video.
To improve content harmonization, we propose a different approach: repeat the sampling process with AnchorExtAttn (Eq. 3) multiple times, progressively reducing the level of noise added at each step. This iterative approach gradually refines the new content’s interaction with the original scene. As shown in Fig.3, we update $\bm{x}_{res}$ representing the difference between the generated output $\mathcal{\hat{V}}_{comp}$ and the original video $\mathcal{{V}}_{orig}$ within the regions where new content appears, allowing each iteration to adjust the generated content’s high-frequency details to better match the original video. We summarize our method in Alg. 1.
Algorithm 1 DynVFX Algorithm
Input:
$\mathcal{V}_{\text{orig}}$, $\mathcal{P}_{\text{VFX}}$triangleright Input video & instruction prompt
$\mathcal{\tau}_{\text{A}}$triangleright Extended Attention threshold
${\Psi}$triangleright Video segmentation model
VLM $\triangleright$ Vision Language model
Preprocess:
$\mathcal{P}_{\text{comp}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Composition Prompt
$\mathcal{O}_{\text{orig}},\mathcal{O}_{\text{edit}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Original objects and VFX object
${M}_{orig}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{V}_{\text{orig}},\mathcal{O}_{\text{orig}})$triangleright Extract source masks
${x}_{orig}\leftarrow\texttt{Encode}[\mathcal{V}_{\text{orig}}]$triangleright Encode video into latent space
$\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}\leftarrow\text{DDIM-Inv}[{x}_{orig}]\quad\forall t\in[T]$
For $t=\tilde{T},\dots,T_{min}$ do
$\bm{x}_{res}=\bm{0}$triangleright initialize the residual latent
$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
$\mathbb{\text{if}}\ t>\mathcal{\tau}_{A}\ \mathbb{\text{then}}\ {K^{E},V^{E}}\leftarrow\mathcal{F}(K_{\text{orig}},M_{\text{orig}}),\mathcal{F}(V_{\text{orig}},M_{\text{orig}})$
else $\mathbb{\text{else }}{K^{E},V^{E}}\leftarrow\emptyset$
$\bm{\hat{x}}_{comp}\leftarrow\texttt{Sampling}[{x}_{comp},\mathcal{P}_{\text{comp}},t;\texttt{AnchorExtAttn}{[K^{E},V^{E}]}]$
$\mathcal{\hat{V}}_{comp}\leftarrow\texttt{Decode}(\bm{\hat{x}}_{comp})$triangleright Decode latent
$\bm{M}_{VFX}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{\hat{V}}_{comp},\mathcal{O}_{\text{edit}})$triangleright Extract VFX masks
$\bm{x}_{res}=\bm{M}_{VFX}\cdot({\hat{x}}_{comp}-{x}_{orig})$
$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
${\mathcal{V}}_{\text{comp}}\leftarrow\texttt{Decode}[{x}_{comp}]$triangleright Output video
Output: $\mathcal{V}_{\text{comp}}$
Figure 6. Qualitative comparison. Sample results of our method, SDEdit (Meng et al., 2022), DDIM inversion (Song et al., 2020), Lora fine-tuning (Hu et al., 2021), and Gen-3 (R Team, Runway, [n. d.]). See SM for videos.
5. Results
We evaluated our method on a dataset of 18 publicly sourced videos, featuring a wide range of complex scenes in terms of camera and object motion, lighting conditions, and physical environments. Additionally, we use some videos from DAVIS (Pont-Tuset et al., 2017). Our videos and implementation details are available in the SM.
Figures 15 show sample results of our method. As seen, our method facilitates natural integration of broad range of visual effects, ranging from environmental effects (tsunami in Fig. 4 and explosion in Fig. 5) to new object insertion (horse riding knight in Fig. 5 and dancing bear in Fig. 6). In all examples, the new content is naturally localized in the scene, even in challenging scenarios of multiple objects (dinosaurs or workers in Fig. 5) and partial occlusions (puppy in Fig. 1 and giraffe in Fig. 5).
5.1. Qualitative Evaluation
To the best of our knowledge, no existing method has been designed to synthesize dynamic objects in a given real video without user masks or any user information except for the input video itself and a simple instruction. We thus compare our method to the following baselines: (i) SDEdit (Meng et al., 2022) using the same T2V model as ours, (ii) DDIM inversion (Song et al., 2020) and sampling with the target prompt, (iii) LORA fine-tuning (Hu et al., 2021) of the T2V model and sampling with the target prompt and (iv) Gen-3 (R Team, Runway, [n. d.]) video-to-video, designed for video stylization.
Figure 6 shows a qualitative comparison to the baselines. As can be seen, all baselines exhibit different limitations in maintaining scene fidelity while introducing new content. SDEdit (Meng et al., 2022) manages to fulfill the edit prompt, yet the scene has significantly deviated from the original one in terms of appearance, motion, positioning, or scale (e.g., deer in the creek). DDIM inversion is not suitable for editing. LORA fine-tuning suffers from the trade-off between preserving aspects of the original scene and adding new content to the scene. Either over-fitting original scene appearance (e.g., added dragon-dog hybrid), or under-fitting the original layout (e.g., incorrect scale of deer). Gen-3 is conditioned on depth maps extracted from the input video, hence tends to significantly alter scene appearance and not allow the insertion of new objects that change the scene layout. In each case, these limitations affect the overall scene coherence and realism of the added elements. Our method successfully adds new content to the scene, achieving high fidelity to the user instructions while allowing for natural interactions between original and added elements (e.g. natural interaction between woman and bear).
Method Metrics VLM-based evaluation User Study
CLIP Directional SSIM Text Alignment Visual Quality Edit Harmonization Dynamics Score Content Integration Edit Harmonization
Gen-3 0.130 0.285 0.418 0.610 0.374 0.379 97.65 93.33
LORA finetuning 0.277 0.361 0.812 0.787 0.756 0.759 92.22 81.11
DDIM inv. sampling 0.184 0.444 0.535 0.699 0.528 0.529 99.20 98.67
SDEdit (0.9) 0.272 0.332 0.794 0.799 0.754 0.756 98.91 82.13
SDEdit (0.6) 0.111 0.567 0.510 0.704 0.513 0.504 97.69 96.76
w/o AnchorExtAttn 0.317 0.697 0.775 0.724 0.683 0.691 89.30 88.89
w/o Iterative Refinement 0.295 0.760 0.817 0.789 0.769 0.760 85.80 86.42
Ours 0.311 0.775 0.860 0.803 0.796 0.785 - -
Table 1. Quantitative Evaluation. We assess the quality of our method compared to several baselines.
5.2. Quantitative Evaluation
We numerically evaluate our results using the following metrics:
(i) Edit fidelity. Following previous works (e.g., (Hsu et al., 2024; Tewel et al., 2024)), we measure per-frame Directional CLIP Similarity (Radford et al., 2021; Gal et al., 2021) to assess the alignment between the change in the source and the target prompt, and the change between the source and edited frames.
(ii) Original content preservation. We evaluate how well the edited video preserves the original content outside the modified region. To this end, we segment the new content in the edited video using (Zhang et al., 2024) , and compute the masked Structural Similarity Index (SSIM) over the complementary regions to the edited ones.
(iii) VLM quality evaluation. We employ a Vision-Language Model (VLM) for expanding the per-frame metrics above in the following manner. We input to the VLM several frames from the edited videos and instruct it to evaluate four key aspects: how well the edit follows the text prompt (Text Alignment), the overall visual quality of the edited frames (Visual Quality), how well the new content is harmonized with the source frames (Edit Harmonization and the realism of the added object’s dynamics relative to the scene (Dynamics Score). For each aspect, the VLM outputs a score between 0 and 1, with higher scores indicating better performance. Our evaluation protocol is included in the SM.
Figure 7. Metrics. We measure CLIP Directional score (higher is better) and masked SSIM (higher is better). Our method demonstrates a better balance between these two metrics.
Figure  7 and Table 1 present the results of the described metrics on a set of 27 video-edit text pairs comprising 20 unique videos. As shown, our method outperforms the baselines in both SSIM and Directional CLIP metrics, demonstrating superior edit fidelity and maintaining higher structural similarity in the unedited regions. The VLM-based evaluation aligns with this assessment and additionally shows that our method produces videos that achieve better content integration and greater motion realism.
(iv) User study. We conducted a user study to evaluate the ability to integrate the new content while preserving the original video. Participants were shown the input video, a text description of the new content, our result, and a baseline output. They were asked two questions: “Which video better preserves the original footage while adding new content?” and “Which video better integrates the new content in a realistic and seamless way?”. In total, we collected 3240 user judgments from 120 users. As seen in Table 1, our method is consistently preferred over all baselines.
5.3. Ablations
We ablate key design choices of our method: anchor extended attention and iterative updates of the edit, by excluding each component from our framework.
As seen in Fig. 4(c), omitting AnchorExtAttn leads to new content being misaligned relative to the original scene - the added content is poorly integrated into the original scene. Applying only the first iteration of our method (w/o iterative refinement, Fig. 4(d)) results in a better alignment with the input video, but the composition is still unstable, as evident, for example, in the puppy’s body hovering over the box in the first scene. Our full method achieves better composition with proper spatial relationships, demonstrating the importance of both components for realistic scene editing Fig. 4(f). We numerically evaluate each ablation with the same set of metrics described in Sec 5.2 and report them in Table 1.
Figure 8. Limitations. In some cases, the T2V diffusion model struggles to precisely follow the edit prompt
6. Discussion and Conclusions
We introduced the task of augmenting real videos with new dynamic content based on a user-provided instruction. We presented a zero-shot method utilizing the T2V diffusion model in a feature manipulation framework, enabling correct localization and natural blending of new content with existing video elements.
As our method is built upon the pre-trained T2V diffusion model, the quality of the generated edits is inherently tied to the performance and capabilities of the underlying model. As seen in Fig. 8, the T2V model sometimes struggles with generating videos precisely following the edit prompt. Additionally, the text-based localization relies on the capabilities of the segmentation model (Zhang et al., 2024), which can sometimes produce inaccurate masks and fails to account for effects like shadows and reflections if not specified in the text prompt. Despite the limitations, our method significantly improves over baselines, expanding the capabilities of pre-trained text-to-video diffusion models.
7. Acknowledgement
This project received funding from the Israeli Science Foundation (grant 2303/20).
Appendix A Implementation Details
A.1. Models
Text-to-Video Model. We use a publicly available CogVideoX-5B (Hong et al., 2022; Yang et al., 2024) text-to-video model, which can generate videos with up to 480x720 pixel resolution, 6 seconds in length, 49 frames at 8 fps. This model is a transformer-based model that processes both text and video modalities together.
Segmentation Model. To segment the prominent objects in the video and the newly generated content, we utilize EVF-SAM2 (Zhang et al., 2024) - a text-based video segmentation model based on SAM2 (Ravi et al., 2024).
Visual Language Model.. Our vision-language model of choice is GPT-4o (Achiam et al., 2023), which we use through the official OpenAI API.
A.2. Keys and Values Extraction
Following (Tumanyan et al., 2023; Yatim et al., 2023), to obtain T2V diffusion model intermediate latents, we use DDIM inversion (applying DDIM sampling in reverse order) on the input video, using 250 forward steps, with an empty string as text prompt. During the forward pass in our method, the intermediate latents are used for the extraction of keys and values.
A.3. Latent Mask Extraction
As discussed in Sec. 4.3, we iteratively update the residual latent $\bm{x}_{res}$ in the regions where the new content appears. This requires calculating the mask of the new content in the latent space. To do this, we first apply the segmentation model (Zhang et al., 2024) to the current output of SDEdit and get the mask of the new content in RGB space. However, the VAE in the T2V diffusion model involves both spatial and temporal downsampling, making it challenging to directly map RGB pixels to their corresponding latent regions. To address this, we encode the RGB masks through the VAE-Encoder and apply clustering to partition the resulting latents into two groups, effectively producing downsampled masks that align with the latent space representation.
A.4. Runtime
Our method’s two most computationally intensive parts are - DDIM inversion, which takes ~15 minutes, and iterative updates of the edit residual, which takes ~20 minutes. Importantly, DDIM inversion needs to be performed only once per video and can support multiple subsequent edits, making the process more efficient when applying various modifications to the same video content.
Appendix B Baselines comparison details
For LORA fine-tuning baseline, we use the following default hyperparameters: Adam optimizer (Kingma and Ba, 2014), $41e-$ 4 learning rank, LORA rank 128, 800 fine-tuning steps. For comparison with GEN-3 (R Team, Runway, [n. d.]), we utilize the Gen-3 Alpha model via the publicly accessible web-based API, setting the ”Structure Transformation” hyperparameter to 5.
Appendix C Additional comparisons
We perform an additional qualitative comparison to MagicVFX (Guo et al., 2024). As can be seen in Fig. 10, MagicVFX struggles to remain faithful to the original scene and has lower visual quality compared to our method.
Figure 10. Comparison to MagicVFX. The result of MagicVFX the output differs significantly from the original video.
Figure 11. Output example for protocol
Figure 12. VLM instructions used for generating the textual descriptions.
Figure 13. VLM evaluation protocol
Appendix D VLM Prompting
While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene. To overcome this, we ask the model to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly. In this conversation, GPT-4o will ”consult” with a VFX artist about how the new content should be integrated into the scene. Based on their discussion, it will be asked to provide a caption that describes how the added content fits into the scene. This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment. In this prompt, we also ask the VLM to provide a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$. The full prompt for the VLM is shown in Figure  12.
In addition, as discussed in Sec. 5.2 we utilize the VLM for interpretable quality assessment. The full set of instructions for the VLM can be seen in Fig. 13.

### v2

@@
 The position, appearance, and motion of the new content are seamlessly integrated into the original footage while accounting for camera motion, occlusions, and interactions with other dynamic objects in the scene, resulting in a cohesive and realistic output video.
-We achieve this via a zero-shot, training-free framework that harnesses a pre-trained text-to-video diffusion transformer to synthesize the new content and a pre-trained Vision Language Model to envision the augmented scene in detail.
+We achieve this via a zero-shot, training-free framework that harnesses a pre-trained text-to-video diffusion transformer to synthesize the new content and a pre-trained vision-language model to envision the augmented scene in detail.
 Specifically, we introduce a novel inference-based method that manipulates features within the attention mechanism, enabling accurate localization and seamless integration of the new content while preserving the integrity of the original scene.
@@
 Text-to-Video Editing, Diffusion Models
-journal: TOG
-Figure 1. DynVFX augments real-world videos with new dynamic content described via simple user-provided text instruction.
+submissionid: 2542journalyear: 2025conference: SIGGRAPH Asia 2025 Conference Papers; December 15–18, 2025; Hong Kong, Hong Kongbooktitle: SIGGRAPH Asia 2025 Conference Papers (SA Conference Papers ’25), December 15–18, 2025, Hong Kong, Hong Kongdoi: 10.1145/3757377.3764008isbn: 979-8-4007-2137-3/2025/12ccs: Computing methodologies Computer vision
+Figure 1.
+DynVFX augments real-world videos with new dynamic content described by a simple user-provided text instruction (shown in the top row under each input video).
+The framework automatically infers where the synthesized content should appear, how it should move, and how it should harmonize at the pixel level with the scene, without requiring any additional user input.
+The key idea is to selectively extend the attention mechanism in a pre-trained text-to-video diffusion model, enforcing the generation to be content-aware of existing scene elements (anchors) from the original video.
+This allows the model to generate content that naturally interacts with the environment, producing complex and realistic video edits in a fully automated way (bottom row).
 1. Introduction
@@
 For instance, the seamless integration of CGI characters, such as Gollum in The Lord of the Rings or T-Rex in Jurassic Park, has empowered filmmakers to blend fantastical elements with real-world environments, resulting in immersive storytelling.
-Inspired by these capabilities, we pose a new creative task: augmenting real-world videos with newly generated dynamic content.
-Specifically, given an input video and a text prompt describing the desired edit, our goal is to synthesize new dynamic objects or complex scene effects, which naturally interact with the existing scene across time (Fig.
+Professional video editing software tools provide artists with precise control over visual effects integration into video footage (Autodesk, 2024; Blender, 2024; Unreal Engine, 2024; Adobe, 2024).
+However, they require significant expertise, manual effort, pre-made assets, and financial resources.
+To disentangle these inspiring storytelling capabilities from professional and economic status, we propose a new creative task: free-vocabulary, text-driven integration of newly generated dynamic content into real-world videos.
+Specifically, given an input video along with a short user instruction describing the new content (e.g., “add a massive whale”), our goal is to synthesize new dynamic objects or complex scene effects that naturally interact with the existing scene across the entire video (Fig.
 1).
+Importantly, our focus is to achieve this without additional user input, allowing anyone to insert dynamic content into videos just by describing it.
+We approach this fully automated task by inferring where the synthesized content should appear, how it should move, and how it should harmonize at the pixel level with the scene without requiring any additional user input.
+This is achieved by harnessing dynamic generative priors learned by a pre-trained text-to-video diffusion transformer (DiT) model and comprehensive scene understanding learned by a vision-language model without any fine-tuning or additional training.
+This contrasts with previous methods that rely on additional user input, such as VFX assets (Hsu et al., 2025) or per-frame references to specify the placement of the new content over time.
+These methods struggle with integrating complex dynamic content (e.g., synthesizing multiple articulated objects or global effects, as shown in the dinosaurs and tsunami examples in Fig.
+5).
 Our task poses several new fundamental challenges.
-First, the generation must be content-aware, such that the position, appearance, and motion of the synthesized dynamic content integrate naturally with the original scene.
-This entails synthesizing objects that respect occlusions, maintain appropriate relative size and perspective with respect to the camera position, and realistically interact with other dynamic objects.
+First, the generation process of the fixed pre-trained DiT model must be content-aware such that the position, appearance, and motion of the synthesized dynamic content integrates naturally with the original scene.
+This entails synthesizing objects that respect occlusions, maintain appropriate relative size and perspective relative to the camera position, and realistically interact with other dynamic objects.
 All of this must be achieved while maintaining the integrity of the original video, ensuring that new content enhances the scene without compromising its authenticity.
-Our method leverages a pre-trained text-to-video model without any fine-tuning or additional training.
-Specifically, given an input video along with a short user instruction describing the new content (e.g., “add a massive whale”), our method produces an edited video where the new content is seamlessly integrated into the original video.
-We achieve it by estimating the residual to the latent representation of the original video.
-To envision the edited scene and to identify prominent foreground objects and visual elements, we leverage a vision-language model that translates the user’s instructions into detailed prompts for the text-to-video diffusion model.
-As our task requires careful placement of the new content, we propose to steer the localization of the edit through Anchor Extended Attention - incorporate a specific set of keys/values extracted from the original video as additional context to the model.
-Additionally, to improve the edit harmonization with the original scene, we propose to iteratively update the estimated edit residual.
+Existing text-based methods focus on controlling diffusion generation to preserve original scene aspects.
+However, this imposes a trade-off between preserving aspects of the original scene and adding new content to the scene: either over-fitting or under-fitting the original video in terms of appearance, motion, positioning, or scale.
+Attention manipulation within a UNet-based architecture has been widely used for content preservation in video editing tasks.
+Extending attention across frames in these models allows for global appearance coherence.
+Yet, spatiotemporal information is preserved only implicitly.
+Hence, in the context of UNet-based models, this has been used mainly for appearance transfer.
+We revisit extended attention in the context of DiT-based text-to-video models, revealing new insights into their internal representation.
+Unlike UNet-based models, we observe that keys/values in DiTs locally encode corresponding video patches across space and time.
+In particular, applying the same positional embedding to both source and target keys/values during the extended attention operation enables alignment between the generated and original content at corresponding spatiotemporal locations.
+We take advantage of this property to enforce content-aware generation for new content placement.
+Specifically, we steer the localization of the edit through Anchor Extended Attention: we incorporate a specific set of keys/values extracted from the original video as additional context to the DiT-based model during sampling, enabling the model to focus on existing scene elements essential for correctly placing and integrating new content naturally.
+Despite accurate placement from Anchor Extended Attention, it does not guarantee pixel-level alignment with the original scene.
+To achieve better harmonization, we iteratively refine the edit by repeatedly sampling with Anchor Extended Attention while updating only the edited regions in each iteration.
+To allow a fully automated framework, a vision-language model is utilized to interpret the interaction and reason about integration.
+We guide the model to: (1) translate the user’s instructions into detailed scene descriptions and (2) identify both the prominent visual elements of the existing scene and the new content to be introduced.
+These identified objects are used in our framework for the application of a text-based segmentation model to aid the localization and harmonization of the new content.
+We demonstrate the effectiveness of our approach on a variety of edits applied to real-world videos.
+Our method supports a wide range of scene augmentations across various scenarios while maintaining realistic interaction, occlusion, lighting, and camera motion.
 It is worth noting that our method utilizes a publicly available text-to-video model, which exhibits a significant gap in video generation quality compared to recent state-of-the-art video models.
-Nevertheless, we observe that within our problem formulation and objective, we can distill from this model surprisingly powerful generative capabilities.
-We demonstrate the effectiveness of our approach on a variety of edits applied to real-world videos.
+Nevertheless, we observe that within our problem formulation and objectives, we can distill surprisingly powerful generative capabilities from this model.
 To summarize, our work makes the following contributions:
-  • We introduce a new task of integrating newly generated dynamic content into real-world videos without relying on the user to provide complex references of the effect (for example, a VFX asset or masks to specify where to locate the VFX).
-  • We propose a tuning-free, zero-shot method that enables harmonized content integration while maintaining high fidelity to the original scene
-  • We propose an automatic VLM-based evaluation metric tailored for our task, considering multiple factors, including original content preservation, new content harmonization, overall visual quality, and alignment with the edit prompt.
-  • We demonstrate state-of-the-art results compared to competing methods, achieving the best trade-off between synthesizing new dynamic elements and maintaining high fidelity to the original content.
+  • We introduce a new task of integrating newly generated dynamic content into real-world videos without relying on the user to provide complex references for the effect.
+  • We propose a tuning-free, zero-shot, fully automated method that enables harmonized content integration while maintaining high fidelity to the original scene.
+  • We demonstrate state-of-the-art results, achieving the best trade-off between synthesizing new dynamic elements and maintaining high fidelity to the original content.
 2. Related Work
-Text-to-Video Models.
-With the rise of large-scale video-text datasets, there has been significant progress in training text-to-video models using new architectures (Bar-Tal et al., 2024; Polyak et al., 2024).
-While the foundational architecture of diffusion models has been commonly linked to inflating text-to-image U-Net-based models with temporal layers (Wang et al., 2023a; cerspense, 2023; Guo et al., 2023), recently, a new family of Transformer-based models (OpenAI, 2024; Kong et al., 2025; Yang et al., 2024; Zheng et al., 2024; HaCohen et al., 2024), referred to as Diffusion Transformers (DiTs) (Peebles and Xie, 2023), have gained popularity, as DiTs enhance spatial coherence and enable arbitrary aspect ratio and video-length training.
+Leveraging Text-to-Image Models.
+With the advancement of text-to-image (T2I) models, techniques for image manipulation based on such models have evolved rapidly.
+Among these advancements, notable progress has been made in the task of instruction-based image editing.
+Several works (Zhang et al., 2024d; Sheynin et al., 2024; Brooks et al., 2023; Zhang et al., 2023a) have proposed directly fine-tuning generative models on pairs of original and edited images coupled with user-provided instructions.
+Object insertion in images falls under the broader umbrella of instruction-based editing techniques.
+Recent object insertion methods (Canberk et al., 2024; Wasserman et al., 2025; Winter et al., 2024) leverage inpainting models to create paired image datasets, which are then used to fine-tune image editing models.
+However, extending these approaches to videos presents significant challenges.
+In particular, generating large-scale instruction-paired video datasets can be expensive in both time and computational resources, as it requires manual effort to annotate frames.
+This cost and complexity make it challenging to adapt existing image-based methodologies directly to the video domain.
+Several works explore the manipulation of internal feature representation in T2I models for image and video editing tasks.
+In U-Net–based architectures, attention keys and values were shown to encode global appearance information (Cao et al., 2023), which several video editing methods (Qi et al., 2023; Wu et al., 2023; Khachatryan et al., 2023) exploit by extending self-attention across frames to enforce global appearance consistency.
+For localized appearance coherency, MasaCtrl (Cao et al., 2023) and Consistory (Tewel et al., 2024b) mask and selectively replace or extend attention keys and values.
+Add-It (Tewel et al., 2024a) employs global weighted extended attention for object insertion in images.
+FlowEdit (Kulikov et al., 2024) uses pre-trained flow models for inversion-free local edits.
+In contrast to the methods mentioned above that may fail to preserve the original content, we propose applying extended attention only to specific regions of the source scene, allowing generation to focus on essential elements.
+Figure 2.
+Pipeline.
+Top (pre-processing).
+Given an input video $\mathcal{V}_{\text{orig}}$ and instruction $\mathcal{P}_{\text{VFX}}$, we (i) Apply our VLM protocol for instruction interpretation, to yield a comprehensive scene description $\mathcal{P}_{\text{comp}}$, original objects $\mathcal{O}_{\text{orig}}$ and target object $\mathcal{O}_{\text{edit}}$ descriptions.
+(ii) DDIM invert $\mathcal{V}_{\text{orig}}$ to extract spatiotemporal keys/values $\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}]$.
+Bottom (editing).
+We initialize the composed latent with $\bm{x}_{\text{comp}}=\bm{x}_{\text{orig}}$ and iterate over a list of descending noise levels $t=\tilde{T}\!\rightarrow\!T_{\min}$ used for noising $\bm{x}_{\text{comp}}$.
+At each iteration $t$ we: (i) noise $\bm{x}_{\text{comp}}$ to noise level $t$, and sample with Anchored Extended Attention, to output $\bm{\hat{x}_{\text{comp}}}$.
+(ii) Update $\bm{x}_{\text{comp}}$ within the new contents masked regions $\bm{{M}_{\text{VFX}}}$ by adding the residual $\bm{x}_{\text{res}}=\bm{{M}_{\text{VFX}}\cdot(\hat{x}_{\text{comp}}-x_{\text{orig}}})$ to $\bm{x}_{\text{orig}}$.
+Repeating this loop gradually integrates the new content, yielding the edited video $\mathcal{V}_{\text{comp}}$.
+Controllable Video Generation and Editing.
+Recently, numerous methods have been developed to incorporate various forms of control signals into video generation pipelines.
+Several video-to-video methods propose to condition the generation on per-frame spatial maps such as depth maps and edge maps (Wang et al., 2023; Zhang et al., 2024b).
+A common approach in video editing is to enforce the preservation of the original scene layout and motion for various tasks.
+Several consistent video editing works (Bar-Tal et al., 2022; Lee et al., 2023; Geyer et al., 2023; Cong et al., 2023) restrict the preservation of scene structure and motion, and rely on a 2D generative prior.
+Therefore, they cannot support large structural deviations or motion adaptation.
+(Yatim et al., 2024; Zhao et al., 2024; Park et al., 2025; Jeong et al., 2024; Materzyńska et al., 2024) proposed to utilize a text-to-video (T2V) model for the task of motion transfer by controlling diffusion video generation to preserve original scene motion.
+(Ku et al., 2024; Ouyang et al., 2024) adopt a two-stage pipeline: first, apply an off-the-shelf image editing method to the video’s first frame, then use an image-to-video generation model to propagate that edit through time via temporal feature injection.
+All of these methods fail to insert new dynamic content.
+Our approach focuses on integrating additional dynamic elements into the video.
+Reference-Based Video Content Insertion.
+Recent methods (Bar-Tal et al., 2024; Ma et al., 2024; Zhang et al., 2024c; Zi et al., 2025; Zhou et al., 2023; Mou et al., 2024; Tu et al., 2025; Bian et al., 2025) have explored the adaptation of video models for video inpainting by conditioning on a masked video and a corresponding binary mask.
+This setup encourages the model to preserve unmasked information while generating new content in the masked region.
+Some methods, such as ReVideo (Mou et al., 2024) and VideoAnydoor (Tu et al., 2025), additionally guide the motion of the new content based on user-provided motion trajectories and bounding boxes.
+VideoDoodles (Yu et al., 2023) combines hand-drawn animations with video footage in a scene-aware manner by tracking a user-provided planar canvas in 3D.
+However, it requires key-frame edit placement.
+Our task fundamentally differs from these methods, which all require explicit per-frame control over where (via per-frame masks (Bian et al., 2025) or bounding boxes (Mou et al., 2024)) or how the new content will move (via motion trajectories (Mou et al., 2024)).
+In contrast, our method does not constrain either aspect and requires only a simple text instruction as input.
+The locations and dynamics in our framework emerge from the generative prior of the T2V model.
+Furthermore, the reliance of these methods on user-provided per-frame annotations is impractical for integrating complex dynamics (e.g., Fig.
+5 tsunami and dinosaurs).
+While a static object can be masked with a simple bounding box, manually defining per-frame masks for complex motion or interactions is extremely difficult.
+Other approaches (KlingAI, 2024; Pika, 2024) reference the asset itself by requiring the user to additionally input an image of the object.
+This approach limits their applicability whenever a suitable example cannot be found for the given video.
+For instance, these methods struggle to generate global effects, making them less preferable than text-based insertion methods.
+Language Models for Video Content Creation.
+Advancements in Vision-Language Models (VLMs) have enabled methods to utilize such models in various video-related tasks.
+Some methods (Chen et al., 2024; Yang et al., 2024) use VLMs to produce detailed video captions from a series of frames, which are then utilized to train T2V generative models.
+Other methods utilize such models for achieving better generation controllability.
+For instance, VideoDirectorGPT (Lin et al., 2023) utilizes a VLM for multi-scene video generation by training diffusion adapters to incorporate additional conditioning inputs, while LVD (Lian et al., 2023) incorporates layout guidance from the VLM during the sampling process.
+AutoVFX (Hsu et al., 2025) uses an LLM to generate a video editing program pipeline based on the user’s instructions.
+In our work, we employ a VLM as a “VFX assistant”.
+Based on a short user instruction, it yields a comprehensive description of the edited video along with the prominent objects present in the scene.
+3. Preliminaries
+Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are a class of generative models that aim to learn a mapping from noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a data distribution $q$.
+Starting from a Gaussian i.i.d.
+noise sample $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the diffusion model $\Phi$ is applied iteratively through a sequence of denoising steps, ultimately producing a clean output sample $\bm{x}_{0}$.
+Recently, a new class of latent T2V models, built on Diffusion Transformers (DiTs) (Peebles and Xie, 2023), has gained significant popularity, as DiTs enhance spatial coherence and enable training for arbitrary aspect ratios and video lengths.
+Unlike previous U-Net–based architectures, which utilize interleaving 2D spatial attention and 1D temporal attention layers, DiT treats the entire video as a long token sequence, where both text and spatiotemporal tokens are processed with full self-attention layers, a mechanism often referred to as 3D attention.
+In each DiT block, text tokens and spatiotemporal tokens are projected into queries, keys, and values using separate sets of weights for each modality, and the sequences of the two modalities are concatenated as a joint input for the attention operation.
+The attention (Vaswani et al., 2017) operation computes the affinities between the d-dimensional projections $\mathbf{Q}$ and $\mathbf{K}$ to yield the output of the layer:
+(1) $\mathbf{A}\cdot\mathbf{V}\ \text{where}\ \mathbf{A}=\texttt{Attention}(\mathbf{Q},\mathbf{K})=\texttt{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)\,.$
+To capture inter-token relationships, Rotary Position Embeddings (RoPE) (Su et al., 2024) are applied to the input queries and keys in the attention operation.
 In this work, we utilize a publicly available DiT-based model (Yang et al., 2024), CogVideoX, for augmenting real-world videos with newly generated dynamic content in a zero-shot manner.
@@
 (a-b) SDEdit suffers from the original scene preservation/edit fidelity trade-off.
-(c-e) Three Extended Attention variants during sampling demonstrate different control levels: Full Extended Attention closely reconstructs the input scene, Masked Extended Attention proves too constrained in overlapping regions despite allowing new content emergence, and our Anchor Extended Attn.
+(c-e) Three Extended Attention variants during sampling demonstrate different control levels: Full Extended Attention closely reconstructs the input scene, Masked Extended Attention proves too constrained in overlapping regions despite allowing new content emergence, and our Anchor Extended Attention.
 achieves optimal results by applying dropout – extending attention only at sparse points within selected regions.
-Object Insertion in Images.
-With the advancement of text-to-image models, techniques for image manipulation leveraging such models have also evolved rapidly.
-Among these advancements, notable progress has been made in the task of instruction-based image editing.
-Several works (Zhang et al., 2023c; Sheynin et al., 2023; Brooks et al., 2023; Zhang et al., 2023a) have proposed to directly fine-tune generative models on pairs of original and edited images coupled with user-provided instructions.
-EmuEdit (Sheynin et al., 2023) leverages a diffusion model trained on a large synthetic dataset to perform various editing tasks guided by task embeddings.
-The task of object insertion into images belongs to the same category and can be considered a subfield of instruction-based editing methods.
-For instance, EraseDraw (Canberk et al., 2024), Paint By Inpaint (Wasserman et al., 2024), and ObjectDrop (Winter et al., 2024) leverage inpainting models to create paired-image datasets, which are then used to fine-tune image editing models.
-However, extending these approaches to videos presents significant challenges.
-In particular, generating large-scale instruction-paired video datasets can be prohibitively expensive both in time and computational resources, as it requires substantial manual effort to annotate frames and ensure alignment between textual instructions and video content.
-This cost and complexity make it challenging to adapt existing image-based methodologies directly to the video domain.
-Concurrently to our method, Add-It (Tewel et al., 2024) proposes to manipulate the attention features of a pre-trained text-to-image diffusion model to insert objects into images in a training-free manner.
-While their method relies on weighted global extended attention, we propose to apply extended attention only to specific regions of the source scene, allowing the generation to focus on essential elements.
-Controllable Video Generation.
-Recently, numerous methods have been developed to incorporate various forms of control signals into video generation pipelines.
-Several video-to-video methods propose to condition the generation on per-frame spatial maps such as depth maps and edge maps (Wang et al., 2023b; Chen et al., 2023).
-A line of work (Yatim et al., 2023; Zhao et al., 2023; Park et al., 2024; Jeong et al., 2023; Geyer et al., 2023) proposed to utilize a text-to-video model for the task of motion transfer.
-Unlike these methods, which are not designed to deviate from the existing structures within a video, our approach focuses on integrating additional dynamic elements into the video.
-Recent methods (Bar-Tal et al., 2024; Ma et al., 2024; Zhang et al., 2023b; Zi et al., 2024; Zhou et al., 2023) have explored adapting text-to-video models for video inpainting by conditioning on a masked video and a corresponding binary mask.
-This setup encourages the model to preserve unmasked information while generating new content in the masked region.
-All of these methods require user-provided masks — an impractical requirement for integrating complex dynamics as it requires anticipating the placement of dynamic objects (e.g., Fig.
-5 jellyfish, tsunami, dinosaurs) ahead of time.
-In contrast, our method allows for automatic new content localization without any user-provided masks.
-While a static object can be masked with a simple bounding box, manually defining masks for complex motion or interactions per frame is extremely difficult.
-Recently, generative Omnimatte (Lee et al., 2024) proposed a method to automatically decompose a video into object layers and their corresponding effects.
-However, it is not designed for the task of new content generation, as it allows only for the removal of existing scene elements.
-VideoDoodles (Yu et al., 2023) combines hand-drawn animations with video footage in a scene-aware manner by tracking a user-provided planar canvas in 3D.
-However, it does not support the creation of non-planar animations.
-Our framework includes both localizing dynamic new content and integrating in a scene-aware manner without reliance on user-provided positions.
-Language Models for Video Content Creation.
-Advancements in Vision Language Models (VLMs) have enabled methods to utilize such models in various video-related tasks.
-Some methods (Chen et al., 2024; Yang et al., 2024) use VLMs to produce detailed video captions from a series of frames, which are then utilized to train text-to-video generative models.
-Other methods utilize such models for achieving better generation controllability.
-For instance, VideoDirectorGPT (Lin et al., 2024) utilizes a VLM for multi-scene video generation by training diffusion adapters to incorporate additional conditioning inputs, while LVD (Lian et al., 2023) incorporates layout guidance from the VLM during the sampling process.
-AutoVFX (Hsu et al., 2024) uses an LLM to generate a video editing program pipeline based on the user instruction.
-In our work, we employ a VLM as a “VFX assistant” that, based on a short user instruction, provides a comprehensive description of the edited video along with the prominent objects present in the scene.
-Professional Software For Video Animation In professional visual effects production, tools such as Autodesk Maya (Autodesk, INC., [n.
-d.]), Blender (Community, 2018), Unreal Engine (Epic Games, [n.
-d.]), Adobe After Effects (Christiansen, 2013) and Houdini (SideFX Houdini FX.
-SideFX, [n.
-d.]) are widely used for creating and compositing complex visual effects.
-These tools provide artists with precise control over object modeling, animation, and integration into video footage.
-While powerful, they require significant expertise, extensive manual intervention, and detailed inputs, such as 3D scene reconstruction or motion tracking.
-All the aforementioned software count on the user to provide the 3D assets.
-Even though creating input 3D assets has become an easier task to solve, using new datasets (Deitke et al., 2023; Qiu et al., 2024), or generating assets based on a user prompt (For example, Meshy.ai and Alpha3D), still, 3D physical elements (like fluid or explosion) or global/multi-object effects present a significant challenge.
-In this work, we take the first steps towards a novice user-friendly workflow.
-Figure 3.
-DynVFX pipeline.
-Top row: Given an input video $\mathcal{V}_{\text{orig}}$, we apply DDIM inversion (see Sec.
-3) and extract spatiotemporal keys and values $[\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}]$ from the original noisy latents.
-Given the user instruction $\mathcal{P}_{\text{VFX}}$ we instruct the VLM to envision the augmented scene and output the text edit prompt $\mathcal{P}_{\text{comp}}$, prominent object descriptions $\mathcal{O}_{\text{orig}}$ that are used to mask out the extracted keys and values and target object descriptions $\mathcal{O}_{\text{edit}}$.
-Bottom row: We estimate a residual $\bm{x}_{\text{res}}$ to the original video latent ($\bm{x}_{\text{orig}}$).
-This is done by iteratively applying SDEdit with our Anchor Extended Attention, segmenting the target objects ($\mathcal{O}_{\text{edit}}$) from the clean result, and updating $\bm{x}_{\text{res}}$ accordingly.
-3. Preliminaries
-Diffusion probabilistic models (Ho et al., 2020; Sohl-Dickstein et al., 2015) are a class of generative models that aim to learn a mapping from noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$ to a data distribution $q$.
-Starting from a Gaussian i.i.d noise $\bm{x}_{T}\sim\mathcal{N}(0,I)$, the diffusion model $\Phi$ is applied iteratively through a sequence of denoising steps, ultimately producing a clean output sample $\bm{x}_{0}$.
-Recently, a new class of latent text-to-video (T2V) models, built on Diffusion Transformers (DiTs) (Peebles and Xie, 2023), has gained significant popularity.
-These models comprise multi-modal blocks (MMDiT (Esser et al., 2024)) that allow for joint processing of both text and image modalities, allowing each to inform and refine the other’s representations.
-To process both modalities together, first, a pre-trained encoder compresses an RGB video $\mathcal{V}$ both spatially and temporally to a latent space.
-Then, the latent is patchified, and the resulting tokens are concatenated to the text tokens produced by the text encoder (Raffel et al., 2023).
-In each MMDiT block, text tokens and spatiotemporal tokens are projected into queries, keys and values using separate sets of weights for each modality, and the sequences of the two modalities are concatenated as a joint input for the attention operation: $\mathbf{Q}=[\mathbf{Q}_{\text{text}},\mathbf{Q}_{\text{spatio}}]$, $\mathbf{K}=[\mathbf{K}_{\text{text}},\mathbf{K}_{\text{spatio}}]$ and $\mathbf{V}=[\mathbf{V}_{\text{text}},\mathbf{V}_{\text{spatio}}]$.
-The attention (Vaswani et al., 2017) operation computes the affinities between the d-dimensional projections $\mathbf{Q},\mathbf{K}$ to yield the output of the layer:
-(1) $\mathbf{A}\cdot\mathbf{V}\ \text{where}\ \mathbf{A}=\texttt{Attention}(\mathbf{Q},\mathbf{K})=\texttt{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\right)$
-To capture inter-token relationships, Rotary Position Embeddings (RoPE) (Su et al., 2021) are applied to the input to the attention operation.
 4. Method
-Given an input video $\mathcal{V}_{\text{orig}}$ and a textual instruction $\mathcal{P}_{\text{VFX}}$, our goal is to synthesize a new video $\mathcal{V}_{\text{VFX}}$, in which new dynamic elements are seamlessly integrated to the existing scene.
-We address this task by estimating a residual in the latent space of the text-to-video diffusion model, which is added to the latent of the original video.
-The final edited video is obtained by decoding their sum with the text-to-video diffusion model’s VAE decoder
-Tackling this task requires ensuring that the generated content, such as new objects or effects, adheres to the dynamics of the existing scene.
-The location and size of the new content must align with the camera motion and the environment, while its actions and movements must appropriately respond to other dynamic objects present in the scene.
-Our framework (illustrated in Fig.
-3) addresses these challenges by incorporating the following key components:
+Given an input video $\mathcal{V}_{\text{orig}}$ and a textual instruction $\mathcal{P}_{\text{VFX}}$, our goal is to synthesize a new video $\mathcal{V}_{\text{VFX}}$ in which new dynamic elements are seamlessly integrated into the existing scene.
+Tackling this task requires ensuring that the location and size of the new content align with the camera motion and the environment, while its actions and movements must respond appropriately to other dynamic objects present in the scene.
+Our framework (Fig.
+2) addresses these challenges by incorporating the following key components:
   1. (1)
     VLM as a VFX Assistant.
-We utilize a pre-trained VLM to interpret the user’s instructions, reason about the interactions with the scene’s dynamics, and provide a descriptive prompt for the T2V diffusion model by guiding it to act as ”VFX assistant” via a system prompt containing guidelines in the context of our tasks.
+We utilize a pre-trained VLM to interpret the user’s instructions, reason about the interactions with the scene’s dynamics, and identify both the prominent existing elements in the scene and the new content to be added.
+The VLM provides a descriptive prompt of the desired scene augmentation, a list of prominent existing elements, and the new content to be added.
+We achieve this by guiding the VLM to act as a “VFX assistant” via a system prompt containing guidelines in the context of our tasks.
   2. (2)
     Localization via Anchor Extended Attention.
-To steer the localization of the edit and make it content-aware, we propose to utilize Extended Attention during sampling from the T2V DiT model to a set of keys and values extracted from sparse locations in the original video.
+Our approach achieves precise placement of new content by guiding the T2V generation model during sampling to be content-aware.
+We steer the localization of the edit by applying Extended Attention with a sparse set of spatiotemporal locations extracted from the original scene, allowing the generation to focus on essential elements.
+Our approach is based on our observation that in the pre-trained T2V DiT model, keys and values locally encode corresponding video patches in both space and time.
   3. (3)
     Content Harmonization via Iterative Refinement.
-To improve the blending of the generated content with the input video and achieve better harmonization, we iteratively update the estimated residual latent by repeating the sampling process with AnchorExtAttn multiple times, progressively reducing the level of noise added at each step.
+To guarantee precise pixel-level alignment of the generated content with the input video and achieve better harmonization, we iteratively refine the estimated edit by repeating the sampling process with AnchorExtAttn multiple times, while preserving all unedited regions.
+Our method does not require user-provided masks to pre-determine the location of the new content, but rather infers it automatically.
+We leverage a text-based segmentation model (Zhang et al., 2024a) for two purposes: (i) Localization, by segmenting prominent elements from the original scene to apply AnchorExtAttn, which allows the T2V model to infer the natural location and dynamics of the new content based on the original scene context.
+(ii) Harmonization, by segmenting the newly added content to ensure pixel-level alignment with the original scene.
 4.1. VLM as a VFX Assistant
-To create a fully automatic framework requiring only a simple user-provided instruction describing the desired content, we incorporate a Vision-Language Model (VLM) into our framework.
-Specifically, given the user instruction $\mathcal{P}_{\text{VFX}}$, along with evenly spaced keyframes from the original video $\mathcal{V}_{\text{orig}}$, we instruct the VLM (Achiam et al., 2023) to provide a detailed caption - a composition prompt $\mathcal{P}_{\text{comp}}$ describing the new composited scene.
-While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene.
-To overcome this, we guide the model to reply in an in-context matter by asking it to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly.
-In this conversation, VLM will “consult” with a VFX artist about how the new content should be integrated into the scene.
-Based on their discussion, it provides a caption that describes how the added content fits into the scene.
-This results in text prompts that encourage the generated output video to include a natural interaction between the new content and the original environment.
-The guidelines include focusing on (1) spatial and dynamic awareness of existing scene elements, (2) preservation of original scene behaviors, and (3) atmospheric coherence between new and existing content.
-See SM for more details about how we use the VLM to reason about the new scene integration with in-context reasoning for VFX.
-We also use the VLM to obtain a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$.
-See more details about prompting and mask processing in SM.
+To create a fully automatic framework that requires only a simple user instruction, we incorporate a VLM into our framework.
+Specifically, given a user instruction $\mathcal{P}_{\text{VFX}}$, along with keyframes from $\mathcal{V}_{\text{orig}}$, we instruct the VLM (Achiam et al., 2023) to provide a composition prompt $\mathcal{P}_{\text{comp}}$ describing the newly augmented scene.
+We use $\mathcal{P}_{\text{VFX}}$ as the edit prompt for the T2V generation process.
+Hence, the edit prompt must include the following details: (i) original scene elements and actions, (ii) new content elements and actions, and (iii) the interaction between them.
+When simply asking the VLM to yield $\mathcal{P}_{\text{VFX}}$, it often fails to produce captions that are both true to the original video and descriptive of the new content well.
+Therefore, using these for T2V generation can result in edits with low fidelity to the instruction (see examples in Sec.
+E.2 and Fig.
+15 in the SM).
+To resolve this issue, we guide the VLM in an in-context manner by instructing it via a system prompt to imagine a conversation with a VFX artist to obtain a caption that would describe the composited scene correctly.
+We observe that this helps the VLM to achieve better scene understanding.
+Specifically, the system prompt includes guidelines to focus on: (1) spatial and dynamic awareness of existing scene elements, (2) preservation of original scene behaviors, and (3) atmospheric coherence between new and existing content.
+Additionally, we utilize the VLM to yield a list of prominent foreground objects in the original video $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt $\mathcal{O}_{\text{edit}}$.
+See additional details in Sec.
+A.3, Sec.
+A.4, and Sec.
+E in the SM.
+In addition, Figs.
+16-19 in the SM include an overview of the protocol, system prompts, and output examples.
 4.2. Localization via Anchor Extended Attention
-A pivotal aspect of our method is the accurate localization of the new content in the existing scene, ensuring alignment with camera motion, occlusions, and scene depth.
-While the composition prompt can describe the desired location, naive noising-denoising with this prompt introduces a trade-off: As shown in Fig.
-2, using SDEdit (Meng et al., 2022) with a high noising timestep fails to retain the original scene, resulting in misaligned new content.
-In contrast, a low noising timestep, as illustrated in Fig.
-2 (b), limits deviations from the original video.
+A pivotal aspect of our method is the accurate placement of the new content.
+While $\mathcal{P}_{\text{comp}}$ can describe the desired location, naive noising-denoising with this prompt introduces a trade-off: As shown in Figs.
+3 (a)-3(b), using SDEdit (Meng et al., 2022) with a high noising timestep fails to retain the original scene, resulting in misaligned new content, whereas a low noising timestep limits deviations from the original video.
+To tackle the localization challenge, we extend the attention module during sampling to include the input video’s corresponding attention features.
+Specifically, we apply DDIM inversion (Song et al., 2020) to the original video $\mathcal{V}_{\text{orig}}$ and extract the spatiotemporal keys and values $K_{\text{orig}},V_{\text{orig}}$ from the attention module of every block in the network and generation timestep $t$.
+These keys and values are then used to extend the attention mechanism during sampling with $\mathcal{P}_{comp}$ to control the localization of the edit.
+When using all keys and values, the extended attention operation can be expressed as:
+(2) $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K_{\text{orig}}],[V_{\text{VFX}},V_{\text{orig}}])\,.$
+We observe that extending attention to the full set of keys and values approximately reconstructs the original video (Fig.
+3(c)).
+This demonstrates that the keys and values of attention layers in the DiT-based model determine more than just appearance information.
+In contrast to UNet-based architectures, keys and values locally encode corresponding video patches.
+We hypothesize that this occurs because the same positional embedding is applied to $K_{\text{orig}}$ as to $K_{\text{VFX}}$.
+Remarkably, applying the same positional embedding to both source and target keys and values enables alignment between the generated and original content at corresponding spatiotemporal locations.
 Figure 4.
@@
 (d) Removing iterative refinement results in poor harmonization.
-Our full method (e) exhibits good localization and harmonization of the edit
-To tackle the localization challenge, we extend the attention module during sampling to include the input video’s corresponding attention features.
-Specifically, we apply DDIM inversion (Song et al., 2020) to the original video $\mathcal{V}_{\text{orig}}$ and extract the spatio-temporal keys and values $K_{\text{orig}},V_{\text{orig}}$ from the attention module of every block in the network and generation timestep $t$.
-These keys and values are then used to extend the attention mechanism during sampling with $\mathcal{P}_{comp}$ to control the localization of the edit.
-When using all keys and values, the extended attention operation can be expressed as:
-(2) $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K_{\text{orig}}],[V_{\text{VFX}},V_{\text{orig}}])$
-Interestingly, extending keys and values provides more than just global information, but each feature locally encodes corresponding patches in the video.
+Our full method (e) exhibits good localization and harmonization of the edit.
+Selective Attention.
+Based on our observation, we propose restricting the extended attention only to specific positions in the source scene.
+Specifically, we use a selection function $\mathcal{F}$ to determine which keys and values are retained.
+Masked Extended Attention.
+A straightforward choice is region‐based masking: identifying the most critical regions for preserving scene coherence and extending attention with keys and values within these masked regions $M_{\text{orig}}$, i.e., $\mathcal{F}(x)=M_{\text{orig}}\circ x$.
+To get $M_{\text{orig}}$, we ask a VLM for $\mathcal{O}_{\text{orig}}$, a list of foreground elements in $\mathcal{V}_{\text{orig}}$ (typically the most spatially prominent elements within a scene), and obtain the corresponding masks using a text-based segmentation model (see Sec.
+A.3 and Sec.
+A.4 in the SM for more details and examples of $M_{\text{orig}}$).
 As shown in Fig.
-2 (c), extending the attention to the full set of keys and values is sufficient to achieve an approximate reconstruction of the original video.
-We hypothesize that this occurs because the same positional embedding is applied to the $K_{\text{orig}}$ as to $K_{\text{VFX}}$, aligning the positional embeddings in spatiotemporal locations.
-This observation provides strong evidence for the significance of the positional embedding.
-Due to this, the setup proves to be overly restrictive in adhering to $\mathcal{P}_{\text{comp}}$, as illustrated in Fig.
-2(c).
-Selective Attention.
-To overcome this limitation, we propose to restrict the extended attention only to specific positions in the source scene.
-Specifically, we use a selection function $\mathcal{F}$ to determine which keys and values are retained.
-An intuitive strategy is to apply region-based attention with Masked Extended Attention by identifying the most critical regions for preserving scene coherence and extending attention with keys and values within these masked regions $M_{\text{orig}}$, i.e., $\mathcal{F}(\mathcal{A})=M_{\text{orig}}\circ\mathcal{A}$.
-To get $M_{\text{orig}}$, we ask a VLM to provide a list of foreground objects in the original video $\mathcal{O}_{\text{orig}}$ (typically the most spatially prominent elements within a scene) and then we obtain corresponding masks using a text-based segmentation model.
-For example, Fig.
-2(d), where the masked region includes the horse, illustrates the effect of this on the generated content.
-As shown, this setup preserves high fidelity to the original scene within the mask $M_{\text{orig}}$ and allows new content to emerge in the unmasked regions.
-However, this approach proves too constrained in overlapping regions. To address this limitation, we propose our Anchor Extended Attention:
-[…]
-[…]
-:=$\mathcal{F}:=\Biggl{\{}\texttt{Drop}_{FG}(M_{\text{orig}})\cup\texttt{Drop}_{BG}(\sim M_{\text{orig}})~{}\circ~{}\mathcal{A}$
-This formulation introduces dropout within the masked regions $M_{\text{orig}}$, generating a sparse set of anchor points that guide the generation while preserving flexibility.
-This formulation introduces a predominantly content-aware selection of anchors from foreground regions, along with a sparser set of background anchors, to achieve robust and spatially coherent integration of new content (e.g.
-the horse to enforce the knewly added knight to align with its motion).
-As demonstrated in Fig.
-2(e), this balanced approach offers sufficient flexibility for creative edits while preserving key spatial cues from the original scene.
-Figure 5. Sample results of our method. See SM for full vide results.
+3(d), this strategy successfully preserves high fidelity to the original scene within $M_{\text{orig}}$ while allowing new content to emerge in unmasked regions.
+However, Fig.
+3(d) reveals two issues: (i) due to the hard constraint in overlapping regions, it is not suitable for handling natural occlusions, and (ii) the lack of guidance in unmasked regions can lead to inaccurate scale or positioning relative to the input scene.
+Anchor Extended Attention.
+To achieve robust and spatially coherent integration of new content, we propose to selectively extend keys/values in a sparse set of foreground anchors and a sparser set of background anchors.
+Introducing Anchor Extended Attention with its formulation:
+(3) $\texttt{AnchorExtAtt}$= $\texttt{Attn}(Q_{\text{VFX}},[K_{\text{VFX}},K^{E}],[V_{\text{VFX}},V^{E}])\,$,
+$\text{s.t. }K^{E}=\mathcal{F}(K_{\text{orig}})~~and~~V^{E}=\mathcal{F}(V_{\text{orig}})\,$,
+$\mathcal{F}(x)=\texttt{Drop}_{\text{FG}}(M_{\text{orig}})\cup\texttt{Drop}_{\text{BG}}(\sim M_{\text{orig}})~\circ~x\,.$
+For each layer, a selection of anchor points is randomly sampled via dropout, in both masked and unmasked regions.
+To steer the localization of the edit, we extend the attention with anchors that correspond to spatiotemporal locations in $\mathcal{V}_{\text{orig}}$, allowing the generation to focus on essential elements (anchors) to achieve robust and spatially coherent integration of new content.
+(e.g., the knight’s motion is aligned with the horse in Fig.
+3(e)).
+This balanced approach offers flexibility for creative edits while preserving key spatial cues from the original scene.
 4.3. Content Harmonization
@@
 As can be seen in Fig.
-2(e), the legs of the horse are not perfectly aligned with the original horse.
-To guarantee pixel-level alignment, a straightforward approach is to extract a mask of the new content $\bm{M}_{VFX}^{rgb}$ from the sampling with AnchorExtAttn (Eq.
-3) output $\mathcal{\hat{V}}_{comp}$.
-More concretely, by applying a text-based segmentation model using the added object description provided by the VLM.
-The mask can then be used to replace the pixels outside it with the corresponding pixels from the original video: $\mathcal{V}{\text{comp}}[\sim\bm{M}_{VFX}^{rgb}]=\mathcal{V}{\text{orig}}[\sim\bm{M}_{VFX}^{rgb}]$.
+3(e), the legs of the horse are not perfectly aligned with the original video.
+To guarantee pixel-level alignment, a straightforward approach is to extract a mask of the new content $\bm{M}_{\text{VFX}}^{\text{rgb}}$ from the sampling with AnchorExtAttn (Eq.
+3) output $\mathcal{\hat{V}}_{\text{comp}}$.
+More concretely, we obtain this mask by applying a text-based segmentation model using the added object description provided by the VLM.
+The mask can then be used to replace the pixels outside it with the corresponding pixels from the original video: $\mathcal{V}{\text{comp}}[\sim\bm{M}_{\text{VFX}}^{\text{rgb}}]=\mathcal{V}{\text{orig}}[\sim\bm{M}_{\text{VFX}}^{\text{rgb}}]$.
 While this preserves the unaffected regions, it often results in poor harmonization with the input video.
-To improve content harmonization, we propose a different approach: repeat the sampling process with AnchorExtAttn (Eq.
+To improve content harmonization, we propose repeating the sampling process with AnchorExtAttn (Eq.
 3) multiple times, progressively reducing the level of noise added at each step.
 This iterative approach gradually refines the new content’s interaction with the original scene.
-As shown in Fig.3, we update $\bm{x}_{res}$ representing the difference between the generated output $\mathcal{\hat{V}}_{comp}$ and the original video $\mathcal{{V}}_{orig}$ within the regions where new content appears, allowing each iteration to adjust the generated content’s high-frequency details to better match the original video.
-We summarize our method in Alg.
-1.
-Algorithm 1 DynVFX Algorithm
-Input:
-$\mathcal{V}_{\text{orig}}$, $\mathcal{P}_{\text{VFX}}$triangleright Input video & instruction prompt
-$\mathcal{\tau}_{\text{A}}$triangleright Extended Attention threshold
-${\Psi}$triangleright Video segmentation model
-VLM $\triangleright$ Vision Language model
-Preprocess:
-$\mathcal{P}_{\text{comp}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Composition Prompt
-$\mathcal{O}_{\text{orig}},\mathcal{O}_{\text{edit}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$triangleright Original objects and VFX object
-${M}_{orig}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{V}_{\text{orig}},\mathcal{O}_{\text{orig}})$triangleright Extract source masks
-${x}_{orig}\leftarrow\texttt{Encode}[\mathcal{V}_{\text{orig}}]$triangleright Encode video into latent space
-$\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}\leftarrow\text{DDIM-Inv}[{x}_{orig}]\quad\forall t\in[T]$
-For $t=\tilde{T},\dots,T_{min}$ do
-$\bm{x}_{res}=\bm{0}$triangleright initialize the residual latent
-$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
-$\mathbb{\text{if}}\ t>\mathcal{\tau}_{A}\ \mathbb{\text{then}}\ {K^{E},V^{E}}\leftarrow\mathcal{F}(K_{\text{orig}},M_{\text{orig}}),\mathcal{F}(V_{\text{orig}},M_{\text{orig}})$
-else $\mathbb{\text{else }}{K^{E},V^{E}}\leftarrow\emptyset$
-$\bm{\hat{x}}_{comp}\leftarrow\texttt{Sampling}[{x}_{comp},\mathcal{P}_{\text{comp}},t;\texttt{AnchorExtAttn}{[K^{E},V^{E}]}]$
-$\mathcal{\hat{V}}_{comp}\leftarrow\texttt{Decode}(\bm{\hat{x}}_{comp})$triangleright Decode latent
-$\bm{M}_{VFX}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{\hat{V}}_{comp},\mathcal{O}_{\text{edit}})$triangleright Extract VFX masks
-$\bm{x}_{res}=\bm{M}_{VFX}\cdot({\hat{x}}_{comp}-{x}_{orig})$
-$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
-${\mathcal{V}}_{\text{comp}}\leftarrow\texttt{Decode}[{x}_{comp}]$triangleright Output video
-Output: $\mathcal{V}_{\text{comp}}$
+As shown in Fig.
+2, we update $\bm{x_{\text{comp}}}$ by adding a residual latent $\bm{x}_{\text{res}}=\bm{{M}_{\text{VFX}}\cdot(\hat{x}_{\text{comp}}-x_{\text{orig}}})$ to $\bm{x_{\text{orig}}}$, where $\bm{{M}_{{\text{VFX}}}}$ is the latent representation of $\bm{{M}_{\text{VFX}}^{\text{rgb}}}$ (see Sec.
+A.4 in SM for details on computing the latent $\bm{{M}_{\text{VFX}}}$).
+Notably, the addition of $\bm{x}_{\text{res}}$ is equivalent to directly replacing values of $\bm{x_{\text{orig}}}$ with those of $\bm{\hat{x}_{\text{comp}}}$ within $\bm{{M}_{\text{VFX}}}$.
+The final edited video is obtained by decoding their sum with the T2V model’s VAE decoder.
+This allows each iteration to adjust the generated content’s high-frequency details to better match the original video.
+We provide implementation details and summarize our method in Sec.
+A and Alg.
+1 in the SM.
 Figure 5.
-Qualitative comparison.
-Sample results of our method, SDEdit (Meng et al., 2022), DDIM inversion (Song et al., 2020), Lora fine-tuning (Hu et al., 2021), and Gen-3 (R Team, Runway, [n.
-d.]).
-See SM for videos.
+Sample Results of DynVFX.
+Our method supports a wide range of scene augmentations across diverse scenarios while maintaining realistic interaction, occlusion, lighting, and camera motion, for example: a golden retriever consistent with camera movement, transparent wings revealing the woman’s silhouette at sunset, and a tsunami flooding the city yet realistically respecting the car dashboard.
+See SM for full videos.
+Figure 6.
+Qualitative Comparison of Text-Based Methods.
+Sample results comparing our method to SDEdit (Meng et al., 2022), DDIM inversion (Song et al., 2020), Lora fine-tuning (Hu et al., 2022), Gen-3 (Runway, 2024) and FlowEdit (Kulikov et al., 2024).
+As can be seen, our method better augments the original scene with new dynamic content that interacts naturally with existing elements in the scene.
+See SM for full video comparison.
 5. Results
-We evaluated our method on a dataset of 18 publicly sourced videos, featuring a wide range of complex scenes in terms of camera and object motion, lighting conditions, and physical environments.
-Additionally, we use some videos from DAVIS (Pont-Tuset et al., 2017).
+We evaluated our method over a set of 57 video-text edit pairs containing 36 unique publicly available videos from the web and the DAVIS dataset (Pont-Tuset et al., 2017).
+These videos feature a wide range of complex scenes in terms of camera and object motion, lighting conditions, and physical environments.
 Our videos and implementation details are available in the SM.
-Figures 15 show sample results of our method.
-As seen, our method facilitates natural integration of broad range of visual effects, ranging from environmental effects (tsunami in Fig.
-4 and explosion in Fig.
-5) to new object insertion (horse riding knight in Fig.
+Figs.
+1 and  5 show sample results of our method.
+As seen, our method facilitates natural integration of a broad range of visual effects, ranging from environmental effects (a tsunami in Fig.
+4 and an explosion in Fig.
+5) to new object insertion (knight riding a horse in Fig.
 5 and dancing bear in Fig.
 6).
-In all examples, the new content is naturally localized in the scene, even in challenging scenarios of multiple objects (dinosaurs or workers in Fig.
-5) and partial occlusions (puppy in Fig.
-1 and giraffe in Fig.
+In all examples, the new content is naturally localized in the scene, even in challenging scenarios involving multiple objects (dinosaurs or workers in Fig.
+5) and partial occlusions (a puppy in Fig.
+1 and a giraffe in Fig.
 5).
+Figure 7. Comparison to Two-Stage Editing Methods. Left pane: our results. Right pane: AnyV2V and I2VEdit. See SM for video comparisons.
 5.1. Qualitative Evaluation
-To the best of our knowledge, no existing method has been designed to synthesize dynamic objects in a given real video without user masks or any user information except for the input video itself and a simple instruction.
-We thus compare our method to the following baselines: (i) SDEdit (Meng et al., 2022) using the same T2V model as ours, (ii) DDIM inversion (Song et al., 2020) and sampling with the target prompt, (iii) LORA fine-tuning (Hu et al., 2021) of the T2V model and sampling with the target prompt and (iv) Gen-3 (R Team, Runway, [n.
-d.]) video-to-video, designed for video stylization.
-Figure 6 shows a qualitative comparison to the baselines.
+We compare our method to the following baselines: (i) SDEdit (Meng et al., 2022) using the same T2V model as ours, (ii) DDIM inversion (Song et al., 2020) and sampling with the target prompt, (iii) LORA fine-tuning (Hu et al., 2022) of the T2V model and sampling with the target prompt, (iv) Gen-3 (Runway, 2024) video-to-video model, designed for video stylization, (v) FlowEdit (Kulikov et al., 2024) and two-stage editing approaches: (vi) AnyV2V (Ku et al., 2024) and (vii) I2VEdit (Ouyang et al., 2024).
+Figs.
+6 and 7 show a qualitative comparison with the baselines.
 As can be seen, all baselines exhibit different limitations in maintaining scene fidelity while introducing new content.
-SDEdit (Meng et al., 2022) manages to fulfill the edit prompt, yet the scene has significantly deviated from the original one in terms of appearance, motion, positioning, or scale (e.g., deer in the creek).
+SDEdit (Meng et al., 2022) and FlowEdit (Kulikov et al., 2024) both manage to fulfill the edit prompt, yet the scene may significantly deviate from the original scene in terms of appearance, motion, positioning, or scale (e.g., deer in the creek).
 DDIM inversion is not suitable for editing.
-LORA fine-tuning suffers from the trade-off between preserving aspects of the original scene and adding new content to the scene.
-Either over-fitting original scene appearance (e.g., added dragon-dog hybrid), or under-fitting the original layout (e.g., incorrect scale of deer).
-Gen-3 is conditioned on depth maps extracted from the input video, hence tends to significantly alter scene appearance and not allow the insertion of new objects that change the scene layout.
+LORA fine-tuning suffers from the trade-off between preserving aspects of the original scene and adding new content to the scene: either overfitting original scene appearance (e.g., adding a dragon-dog hybrid), or underfitting the original layout (e.g., incorrect scale of deer).
+Gen-3 is conditioned on structure and content representation extracted from the input video, hence it tends to significantly alter scene appearance and does not allow the insertion of new objects that change the scene layout.
+Both AnyV2V and I2VEdit take the first edited frame as additional input, but still struggle to maintain the coherent motion of the generated objects.
+In particular, AnyV2V often fails to accurately propagate the edits through the video, leading to distorted geometries.
 In each case, these limitations affect the overall scene coherence and realism of the added elements.
-Our method successfully adds new content to the scene, achieving high fidelity to the user instructions while allowing for natural interactions between original and added elements (e.g.
-natural interaction between woman and bear).
+Our method successfully adds new content to the scene, achieving high fidelity to the user instructions while allowing for natural interactions between original and added elements (e.g., natural interaction between woman and bear).
+We provide details about the baselines comparison in Sec.
+B in the SM.
+Additionally, we provide qualitative comparisons with VFX reference-based methods (Pika, 2024; KlingAI, 2024) in Sec.
+D and Fig.
+12 in the SM.
+Table 1.
+Quantitative Comparison.
+We compare our method with other video editing methods and ablations using metrics for edit fidelity (CLIP Directional) and original content preservation (masked SSIM), a VLM-based evaluation of text alignment, visual quality, harmonization, and motion realism, and a user study on content integration and edit harmonization.
+User study values indicate the percentage of participants who preferred our method over each baseline.
 Method Metrics VLM-based evaluation User Study
 CLIP Directional SSIM Text Alignment Visual Quality Edit Harmonization Dynamics Score Content Integration Edit Harmonization
-Gen-3 0.130 0.285 0.418 0.610 0.374 0.379 97.65 93.33
-LORA finetuning 0.277 0.361 0.812 0.787 0.756 0.759 92.22 81.11
-DDIM inv. sampling 0.184 0.444 0.535 0.699 0.528 0.529 99.20 98.67
-SDEdit (0.9) 0.272 0.332 0.794 0.799 0.754 0.756 98.91 82.13
-SDEdit (0.6) 0.111 0.567 0.510 0.704 0.513 0.504 97.69 96.76
-w/o AnchorExtAttn 0.317 0.697 0.775 0.724 0.683 0.691 89.30 88.89
-w/o Iterative Refinement 0.295 0.760 0.817 0.789 0.769 0.760 85.80 86.42
-Ours 0.311 0.775 0.860 0.803 0.796 0.785 - -
-Table 1. Quantitative Evaluation. We assess the quality of our method compared to several baselines.
+Gen-3 0.142 0.283 0.410 0.586 0.363 0.374 98.687 94.691
+LORA finetuning 0.292 0.349 0.781 0.755 0.726 0.714 92.221 71.968
+DDIM inv. sampling 0.193 0.440 0.465 0.609 0.448 0.448 98.951 93.145
+SDEdit (0.9) 0.290 0.321 0.778 0.750 0.717 0.730 99.154 75.645
+SDEdit (0.6) 0.105 0.558 0.381 0.703 0.406 0.390 92.655 92.470
+FlowEdit 0.288 0.414 0.771 0.765 0.712 0.736 97.346 65.494
+AnyV2V 0.300 0.468 0.678 0.669 0.615 0.619 91.142 77.747
+I2VEdit 0.288 0.524 0.769 0.773 0.724 0.719 91.030 82.563
+w/o AnchorExtAttn 0.325 0.691 0.765 0.713 0.670 0.687 81.417 81.109
+w/o Iterative Refinement 0.291 0.736 0.749 0.741 0.697 0.696 80.122 80.492
+w/o VLM Protocol 0.254 0.741 0.705 0.755 0.669 0.667 71.360 73.708
+Ours 0.307 0.749 0.843 0.784 0.778 0.775 - -
 5.2. Quantitative Evaluation
@@
 (i) Edit fidelity.
-Following previous works (e.g., (Hsu et al., 2024; Tewel et al., 2024)), we measure per-frame Directional CLIP Similarity (Radford et al., 2021; Gal et al., 2021) to assess the alignment between the change in the source and the target prompt, and the change between the source and edited frames.
+Following previous works (e.g., (Hsu et al., 2025; Tewel et al., 2024a)), we measure per-frame Directional CLIP Similarity (Radford et al., 2021; Gal et al., 2022) to assess the alignment between the change in the source and the target prompt, and the change between the source and edited frames.
 (ii) Original content preservation.
 We evaluate how well the edited video preserves the original content outside the modified region.
-To this end, we segment the new content in the edited video using (Zhang et al., 2024) , and compute the masked Structural Similarity Index (SSIM) over the complementary regions to the edited ones.
+To this end, we segment the new content in the edited video using (Zhang et al., 2024a), and compute the masked Structural Similarity Index (SSIM) over the regions complementary to the edited ones.
 (iii) VLM quality evaluation.
-We employ a Vision-Language Model (VLM) for expanding the per-frame metrics above in the following manner.
-We input to the VLM several frames from the edited videos and instruct it to evaluate four key aspects: how well the edit follows the text prompt (Text Alignment), the overall visual quality of the edited frames (Visual Quality), how well the new content is harmonized with the source frames (Edit Harmonization and the realism of the added object’s dynamics relative to the scene (Dynamics Score).
+Inspired by (Hsu et al., 2025; Huang et al., 2024), we employ a VLM to expand the per-frame metrics above as follows.
+We input several frames from the edited videos into the VLM and instruct it to evaluate four key aspects: how well the edit follows the text prompt (Text Alignment), the overall visual quality of the edited frames (Visual Quality), how well the new content is harmonized with the source frames (Edit Harmonization), and the realism of the added object’s dynamics relative to the scene (Dynamics Score).
 For each aspect, the VLM outputs a score between 0 and 1, with higher scores indicating better performance.
-Our evaluation protocol is included in the SM.
-Figure 7. Metrics. We measure CLIP Directional score (higher is better) and masked SSIM (higher is better). Our method demonstrates a better balance between these two metrics.
-Figure  7 and Table 1 present the results of the described metrics on a set of 27 video-edit text pairs comprising 20 unique videos.
-As shown, our method outperforms the baselines in both SSIM and Directional CLIP metrics, demonstrating superior edit fidelity and maintaining higher structural similarity in the unedited regions.
-The VLM-based evaluation aligns with this assessment and additionally shows that our method produces videos that achieve better content integration and greater motion realism.
+In the SM, we include our evaluation protocol and the justification of its use as an evaluation metric, showing its alignment with human preferences and consistent outcomes with an alternative VLM backbone.
+We include this in Sec.
+F and Table 2 in the SM.
+Table 1 and Fig.
+9 in the SM present the results of the described metrics on a set of 57 video-text edit pairs comprising 34 unique videos.
+As shown, our method outperforms the baselines in both SSIM and Directional CLIP metrics, demonstrating superior edit fidelity while maintaining higher structural similarity in the unedited regions.
+The VLM-based evaluation aligns with this assessment and further shows that our method produces videos that achieve better content integration and greater motion realism.
 (iv) User study.
-We conducted a user study to evaluate the ability to integrate the new content while preserving the original video.
+We conducted a user study to evaluate the ability to integrate new content while preserving the original video.
 Participants were shown the input video, a text description of the new content, our result, and a baseline output.
 They were asked two questions: “Which video better preserves the original footage while adding new content?” and “Which video better integrates the new content in a realistic and seamless way?”.
-In total, we collected 3240 user judgments from 120 users.
+In total, we collected 17,160 user judgments from 330 users.
 As seen in Table 1, our method is consistently preferred over all baselines.
 5.3. Ablations
-We ablate key design choices of our method: anchor extended attention and iterative updates of the edit, by excluding each component from our framework.
+We ablate key design choices of our method: Anchor Extended Attention, iterative updates of the edit, and our VLM as a VFX Assistant protocol, by excluding each component from our framework.
+The ablation of our VLM protocol can be found in Sec.
+E.2 and Fig.
+15 in the SM.
 As seen in Fig.
-4(c), omitting AnchorExtAttn leads to new content being misaligned relative to the original scene - the added content is poorly integrated into the original scene.
+4(c), omitting AnchorExtAttn leads to new content being misaligned relative to the original scene, with the added content poorly integrated into the original scene.
 Applying only the first iteration of our method (w/o iterative refinement, Fig.
 4(d)) results in a better alignment with the input video, but the composition is still unstable, as evident, for example, in the puppy’s body hovering over the box in the first scene.
-Our full method achieves better composition with proper spatial relationships, demonstrating the importance of both components for realistic scene editing Fig.
-4(f).
+Our full method achieves better composition with proper spatial relationships, demonstrating the importance of both components for realistic scene editing (Fig.
+4(f)).
 We numerically evaluate each ablation with the same set of metrics described in Sec 5.2 and report them in Table 1.
-Figure 8. Limitations. In some cases, the T2V diffusion model struggles to precisely follow the edit prompt
+An additional ablation example can be found in Fig.
+14 in the SM.
+Figure 8. Limitations. In some cases, the T2V model struggles to precisely follow the edit prompt.
 6. Discussion and Conclusions
 We introduced the task of augmenting real videos with new dynamic content based on a user-provided instruction.
-We presented a zero-shot method utilizing the T2V diffusion model in a feature manipulation framework, enabling correct localization and natural blending of new content with existing video elements.
-As our method is built upon the pre-trained T2V diffusion model, the quality of the generated edits is inherently tied to the performance and capabilities of the underlying model.
-As seen in Fig.
-8, the T2V model sometimes struggles with generating videos precisely following the edit prompt.
-Additionally, the text-based localization relies on the capabilities of the segmentation model (Zhang et al., 2024), which can sometimes produce inaccurate masks and fails to account for effects like shadows and reflections if not specified in the text prompt.
-Despite the limitations, our method significantly improves over baselines, expanding the capabilities of pre-trained text-to-video diffusion models.
-7. Acknowledgement
-This project received funding from the Israeli Science Foundation (grant 2303/20).
+We presented a zero-shot method utilizing the T2V model in a feature manipulation framework, enabling correct localization and natural blending of new content with existing video elements.
+As our method is built upon the pre-trained T2V model, the quality of the generated edits is inherently tied to the performance and capabilities of the underlying model.
+As shown in Fig.
+8, the T2V model sometimes struggles to generate videos that precisely follow the edit prompt; additional failure cases and limitations are detailed in Sec.
+C and Fig.
+11 in the SM.
+While the quality of the edits is bound by the capabilities of the T2V model, our method is model-agnostic and can be re-implemented with future transformer-based T2V models.
+We anticipate its performance will improve as more powerful models become available.
+Even under these constraints, our method significantly outperforms baselines, expanding the capabilities of pre-trained text-to-video diffusion models.
+Appendix
 Appendix A Implementation Details
@@
 Text-to-Video Model.
-We use a publicly available CogVideoX-5B (Hong et al., 2022; Yang et al., 2024) text-to-video model, which can generate videos with up to 480x720 pixel resolution, 6 seconds in length, 49 frames at 8 fps.
+We use a publicly available CogVideoX-5B (Hong et al., 2022; Yang et al., 2024) text-to-video model, which can generate videos with up to $480\times 720$ pixels resolution, 6 seconds in length, 49 frames at 8 fps.
 This model is a transformer-based model that processes both text and video modalities together.
 Segmentation Model.
-To segment the prominent objects in the video and the newly generated content, we utilize EVF-SAM2 (Zhang et al., 2024) - a text-based video segmentation model based on SAM2 (Ravi et al., 2024).
-Visual Language Model.. Our vision-language model of choice is GPT-4o (Achiam et al., 2023), which we use through the official OpenAI API.
+To segment the prominent objects in the video and the newly generated content, we utilize EVF-SAM2 (Zhang et al., 2024a) - a text-based video segmentation model based on SAM2 (Ravi et al., 2024).
+Visual Language Model. Our vision-language model of choice is GPT-4o (Achiam et al., 2023), which we use through the official OpenAI API.
+Figure 9.
+Metrics Visualization.
+We measure CLIP Directional score (higher is better) and masked SSIM (higher is better).
+Our method demonstrates a better balance between these two metrics.
+See Sec.
+5.2 for full discussion.
+Algorithm 1 DynVFX Algorithm
+Input:
+$\mathcal{V}_{\text{orig}}$, $\mathcal{P}_{\text{VFX}}$ $\triangleright$ Input video & instruction prompt
+$\mathcal{\tau}_{\text{A}}$ $\triangleright$ Extended Attention threshold
+$\Psi$} $\triangleright$ Video segmentation model
+VLM $\triangleright$ Vision Language model
+Preprocess:
+$\mathcal{P}_{\text{comp}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$ $\triangleright$ Composition Prompt
+$\mathcal{O}_{\text{orig}},\mathcal{O}_{\text{edit}}\leftarrow\text{VLM}[\mathcal{V}_{\text{orig}},\mathcal{P}_{\text{VFX}}]$ $\triangleright$ Original objects and VFX object
+${M}_{orig}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{V}_{\text{orig}},\mathcal{O}_{\text{orig}})$ $\triangleright$ Extract source masks
+${x}_{orig}\leftarrow\texttt{Encode}[\mathcal{V}_{\text{orig}}]$ $\triangleright$ Encode video into latent space
+$\mathbf{K}_{\text{orig}},\mathbf{V}_{\text{orig}}\leftarrow\text{DDIM-Inv}[{x}_{orig}]\quad\forall t\in[T]$
+For $t=\tilde{T},\dots,T_{min}$ do
+$\bm{x}_{res}=\bm{0}$ $\triangleright$ initialize the residual latent
+$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
+$\mathbb{\text{if}}\ t>\mathcal{\tau}_{A}\ \mathbb{\text{then}}\ {K^{E},V^{E}}\leftarrow\mathcal{F}(K_{\text{orig}},M_{\text{orig}}),\mathcal{F}(V_{\text{orig}},M_{\text{orig}})$
+$\mathbb{\text{else }}{K^{E},V^{E}}\leftarrow\emptyset$
+$\bm{\hat{x}}_{comp}\leftarrow\texttt{Sampling}[{x}_{comp},\mathcal{P}_{\text{comp}},t;\texttt{AnchorExtAttn}{[K^{E},V^{E}]}]$
+$\mathcal{\hat{V}}_{comp}\leftarrow\texttt{Decode}(\bm{\hat{x}}_{comp})$ $\triangleright$ Decode latent
+$\bm{M}_{VFX}\leftarrow\texttt{Get-Latent-Mask}(\Psi;\mathcal{\hat{V}}_{comp},\mathcal{O}_{\text{edit}})$ $\triangleright$ Extract VFX masks
+$\bm{x}_{res}=\bm{M}_{VFX}\cdot({\hat{x}}_{comp}-{x}_{orig})$
+$\bm{x}_{comp}={x}_{orig}+{x}_{res}$
+${\mathcal{V}}_{\text{comp}}\leftarrow\texttt{Decode}[{x}_{comp}]$ $\triangleright$ Output video
+Output: $\mathcal{V}_{\text{comp}}$
 A.2. Keys and Values Extraction
-Following (Tumanyan et al., 2023; Yatim et al., 2023), to obtain T2V diffusion model intermediate latents, we use DDIM inversion (applying DDIM sampling in reverse order) on the input video, using 250 forward steps, with an empty string as text prompt.
-During the forward pass in our method, the intermediate latents are used for the extraction of keys and values.
-A.3. Latent Mask Extraction
+Following (Tumanyan et al., 2023; Yatim et al., 2024), to obtain the T2V diffusion model intermediate latents, we apply DDIM inversion (applying DDIM sampling in reverse order) on the input video.
+We perform inversion solely to extract the keys and values for AnchorExtAttn, while the sampling itself is initialized from a noisy $x_{comp}$.
+For the inversion process, we use 250 steps, with an empty string as the text prompt.
+We derive keys and values from the noisy latents produced by the inversion process (not from the reconstruction).
+During the sampling iterations in our method, we use these extracted keys and values for AnchorExtAttn.
+(Eq.
+3) randomly samples a selection of anchor points, which are re-selected in each layer of the model.
+For all our results, we fix the hyperparameters and only adjust the dropout percentage in the AnchorExtAttn: 5% dropout for background and 70% for foreground in regular edits, while global edits use 10% for the foreground.
+A.3. Obtaining $M_{orig}$
+Figure 10.
+Qualitative Examples of $M_{orig}$.
+Each example includes the input video, user instruction, original masks, and corresponding result.
+In the top-right whale edit, our VLM protocol identifies the prominent elements to be coral and divers.
+In the top left pirate ships edit, the prominent elements are the cliffs.
+In the bottom left puppy edit, the woman and the ocean are identified.
+And in the bottom right bear edit, the woman and the couch are chosen.
+To get $M_{\text{orig}}$, we ask a VLM to provide a list of foreground objects in the original video $\mathcal{O}_{\text{orig}}$ (typically the most spatially prominent elements within a scene) and then we obtain corresponding masks using EVF-SAM (Zhang et al., 2024a) - a text-based segmentation model.
+Specifically, given the user instruction, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, our VLM protocol identifies the prominent elements from the original scene using the system prompt in Fig.
+18.
+Next, each element (represented by words) is fed to EVF-SAM to produce masks, and their union forms $M_{orig}$, which is used in our AnchorExtAttn.
+Examples of such masks can be seen in Fig.
+10.
+A.4. Latent Mask Extraction
 As discussed in Sec.
@@
 This requires calculating the mask of the new content in the latent space.
-To do this, we first apply the segmentation model (Zhang et al., 2024) to the current output of SDEdit and get the mask of the new content in RGB space.
+To do this, we first apply the segmentation model (Zhang et al., 2024a) to the current output of SDEdit and get the mask of the new content in RGB space.
 However, the VAE in the T2V diffusion model involves both spatial and temporal downsampling, making it challenging to directly map RGB pixels to their corresponding latent regions.
 To address this, we encode the RGB masks through the VAE-Encoder and apply clustering to partition the resulting latents into two groups, effectively producing downsampled masks that align with the latent space representation.
-A.4. Runtime
-Our method’s two most computationally intensive parts are - DDIM inversion, which takes ~15 minutes, and iterative updates of the edit residual, which takes ~20 minutes.
+A.5. Runtime and Memory Usage
+Our method’s most computationally intensive parts are - DDIM inversion, which takes ~15 minutes, and iterative updates of the edit residual, which takes ~20 minutes, while a single iteration of sampling with AnchoExtAttn takes between 2-5 minutes (depending on the noise level).
 Importantly, DDIM inversion needs to be performed only once per video and can support multiple subsequent edits, making the process more efficient when applying various modifications to the same video content.
+Querying the VLM and EVF-SAM adds negligible runtime (~5 and ~20 seconds correspondingly).
+We run our method on a NVIDIA A100, with 72GB memory in usage.
 Appendix B Baselines Comparison Details
-For LORA fine-tuning baseline, we use the following default hyperparameters: Adam optimizer (Kingma and Ba, 2014), $41e-$ 4 learning rank, LORA rank 128, 800 fine-tuning steps.
-For comparison with GEN-3 (R Team, Runway, [n.
-d.]), we utilize the Gen-3 Alpha model via the publicly accessible web-based API, setting the ”Structure Transformation” hyperparameter to 5.
-Appendix C Additional comparisons
-We perform an additional qualitative comparison to MagicVFX (Guo et al., 2024).
+The Baseline runtimes are (i) SDEdit - ~5 minutes (ii) DDIM inversion + Sampling - ~20 minutes (iii) LORA fine-tuning - ~2 hours, (iv) Gen-3 - ~30 seconds, (v) FlowEdit - ~3 minutes, (vi) AnyV2V - ~3 minutes and (vii) I2VEdit - ~30 minutes.
+In terms of utilized Video Models, baselines (i) SDEdit, (ii) DDIM-inversion+Sampling, and (iii) LORA fine-tuning utilized the same text-to-video model (Yang et al., 2024) as our method, (iv) Gen-3 utilizes a proprietary video-to-video model by Runway (Alpha model via the publicly accessible web-based API), (v) FlowEdit is applied with Hunyuan Video model (Kong et al., 2024), (vi) AnyV2V uses I2VGen-XL model (Zhang et al., 2023b), supporting 15 frames, and (vii) I2VEdit utilizes Stable Video Diffusion model (Blattmann et al., 2023) supporting 27 frames.
+For AnyV2V and I2VEdit, the first frames were edited with Gemini image editing (Gemini 2.0, 2024).
+For LORA fine-tuning baseline, we use the following default hyperparameters: Adam optimizer (Kingma and Ba, 2014), 1e41e-4 learning rank, LORA rank 128, 800 fine-tuning steps.
+For comparison with Gen-3 (Runway, 2024), we set the ”Structure Transformation” hyperparameter to 5.
+For FlowEdit, we use source and target prompts produced from our VLM-VFX protocol, as this method heavily relies on source and target prompting.
+Appendix C Limitations
+Figure 11.
+Failure Case Example.
+When trying to add content that already exists in the scene (e.g., an additional dog), it is best to specify the differences between them in the user instructions.
+As can be seen, when trying to add a Poodle (similar appearance and size to a Havanese), EVF-SAM fails to distinguish between the two.
+However, when trying to add a German Shepherd, the segmentation model manages to make the distinction.
+Figure 12.
+Comparison to Reference-Based Insertion Methods.
+The left plane shows the results of our method, while the right plane shows the results of Pika and Kling.
+As can be seen, these reference-based methods struggle to generate global effects.
+Figure 13.
+Comparison to MagicVFX.
+MagicVFX (second column) deviates significantly from the original video.
+As seen in the third column, our method successfully adds new content while maintaining high fidelity to the original scene.
+Our method relies on the target segmentation provided by a text-based segmentation model (Zhang et al., 2024a).
+While this solution often succeeds for most edits, it can sometimes produce inaccurate masks and fail to account for effects like shadows and reflections if not specified in the text prompt.
+To handle such cases, users can add a specification in text to include these secondary effects.
+Furthermore, the segmentation model may struggle in scenarios where multiple entities with similar visual characteristics are present in the scene.
+As shown in Fig.
+11, segmentation issues can occur when adding content of similar appearance and size.
+However, when the entities are distinguishable enough, the new content is successfully integrated into the scene.
+Furthermore, our framework performs a single-shot edit and does not support interactive, in-context refinements - once an edit is applied, users cannot request follow-up adjustments (e.g., tweaking object placement or appearance) without restarting the process.
+While VFX pipelines typically involve several feedback-driven iterations to perfect a scene, our current design does not support such incremental revisions.
+Future work could integrate an interactive editing loop, potentially via an autoregressive T2V model, to allow users to make successive edits without reinitializing the entire workflow.
+Figure 14.
+Additional Example for Ablations.
+(b) Excluding both AnchorExtAttn and the Iterative refinement process results in significant misalignment with the original scene and poor harmonization (e.g., the size of the puppy relative to the scene and boundary artifacts).
+(c) Omitting AnchorExtAttn leads to incorrect positioning of the new content.
+(d) Removing iterative refinement results in poor harmonization.
+Our full method (e) exhibits good localization and harmonization of the edit.
+Appendix D Additional Qualitative Comparisons
+We perform additional qualitative comparisons to MagicVFX (Guo et al., 2024) and two reference-based methods: Pikadditions (Pika, 2024) and Kling (KlingAI, 2024).
 As can be seen in Fig.
-10, MagicVFX struggles to remain faithful to the original scene and has lower visual quality compared to our method.
-Figure 10. Comparison to MagicVFX. The result of MagicVFX the output differs significantly from the original video.
-Figure 11. Output example for protocol
-Figure 12. VLM instructions used for generating the textual descriptions.
-Figure 13. VLM evaluation protocol
-Appendix D VLM Prompting
-While the model gives an accurate, descriptive source scene caption, in some cases, we observed that it fails to give captions suitable for compositing VFX with the scene.
+13, MagicVFX struggles to remain faithful to the original scene and has lower visual quality compared to our method.
+Furthermore, as illustrated in Fig.
+12, both Pikadditions and Kling struggle to generate global effects.
+As can be seen, the audience members are incorrectly overlaid in front of the dancer.
+Additionally, the dinosaur occludes one of the bikers, and only one dinosaur is generated, although we prompt for a global edit.
+This demonstrates scenarios where text-based insertion is preferable over reference-based insertion.
+Even though such methods do not explicitly require the control of location and dynamics of the new content (as our method), they rely on the reference assets, which limits their applicability whenever a suitable example cannot be found for the given video.
+Appendix E VLM Prompting
+E.1. Our VLM as a VFX Assistant Protocol
+Although the model is capable of providing descriptive captions of the source scene, in some cases, we observed that it fails to provide captions suitable for compositing VFX with the scene.
 To overcome this, we ask the model to imagine a conversation with a visual effects (VFX) artist to obtain a caption that would describe the composited scene correctly.
@@
 In this prompt, we also ask the VLM to provide a list of prominent foreground objects in the original video: $\mathcal{O}_{\text{orig}}$ and the object that will be added according to the edit prompt: $\mathcal{O}_{\text{edit}}$.
-The full prompt for the VLM is shown in Figure  12.
-In addition, as discussed in Sec. 5.2 we utilize the VLM for interpretable quality assessment. The full set of instructions for the VLM can be seen in Fig. 13.
+The full prompt for the VLM is shown in Fig.
+17.
+In addition, as discussed in Sec. 5.2, we utilize the VLM for interpretable quality assessment. The full set of instructions for the VLM can be seen in Fig. 19.
+E.2. Ablation of the VLM as a VFX Assistant Protocol
+Figure 15.
+Ablating VLM Protocol.
+Without the protocol (middle), the VLM often yields an unsuitable prompt for adding the desired content, while our protocol (right), gives $\mathcal{P}_{\text{comp}}$ which allows for scene-aware additions to $\mathcal{O}_{\text{orig}}$ (see jellyfish and parrot behind the window).
+Additionally, as can be seen in Table 1 in the main paper, Table 2 and Fig.
+9 from the SM, the metrics for this ablation are worse than those for our full method.
+To validate the importance of our VLM protocol, we performed an ablation by prompting the VLM with a simplified system prompt, asking it to caption the original video and provide the edit prompt.
+As seen in Fig.
+15, simplifying the system prompt results in misinterpretation of the instruction with respect to the scene or fails to add new content in general.
+Appendix F VLM as Evaluation Metric Validation
+Table 2.
+VLM Quality Evaluation with an alternative backbone (Claude Haiku 3.5).
+We report text alignment, visual quality, edit harmonization, and dynamics score.
+The results closely match those obtained with GPT-4o in the main paper, showing consistent outcomes across all metrics and methods.
+Method Text Align. Visual Quality Edit Harm. Dynamics Score
+Gen-3 0.517 0.499 0.412 0.445
+LORA finetuning 0.807 0.755 0.722 0.719
+DDIM inv. sampling 0.423 0.487 0.384 0.385
+SDEdit (0.9) 0.769 0.734 0.697 0.682
+SDEdit (0.6) 0.420 0.561 0.425 0.405
+FlowEdit 0.670 0.651 0.591 0.605
+AnyV2V 0.693 0.683 0.625 0.625
+I2VEdit 0.769 0.747 0.697 0.698
+w/o AnchorExtAttn 0.656 0.620 0.547 0.556
+w/o Iterative Refinement 0.704 0.672 0.631 0.634
+w/o VLM Protocol 0.697 0.700 0.631 0.644
+Ours 0.820 0.760 0.721 0.721
+To show that our VLM-based evaluation aligns with human preferences, we performed the following assessments: Measured Pearson’s correlation coefficients between our VLM scores (Edit Harmonization and Dynamics) and the user study scores for Edit Harmonization, as both assess content integration quality.
+We observe a strong correlation $r=0.82$ ($p=0.025$) and $r=0.89$ ($p=0.007$).
+We compute Pearson’s correlation between the Text Alignment score and the CLIP Directional score - a metric commonly used to assess text alignment and shown to correlate with human preference (Kim et al., 2025) - and found a strong correlation: r=0.81 (p=0.029).
+These strong correlations suggest that our metric aligns with human judgments in the context of our task.
+We utilize the same VLM (GPT-4o (Achiam et al., 2023)) in our pipeline and evaluation.
+However, this does not introduce bias in our evaluation due to the separation between the internal representations of the T2V model and the VLM’s reasoning.
+During evaluation, the VLM does not “recognize” the video generated by the T2V model as its own, making the assessment unbiased and independent of its earlier role.
+We further validate this by evaluating the results with a different VLM (Claude Haiku 3.5 (Anthropic, 2024)) in Table 2 and demonstrate that it provides similar scoring to that reported in the paper.
+Our VLM-based evaluation metrics are inspired by recent work such as AutoVFX (Hsu et al., 2025) and (Wu et al., 2024), which demonstrate the effectiveness of VLMs for automatic quality assessment tasks.
+Figure 16.
+VLM as a VFX Assistant Overview.
+Given instructions and an input video, the VLM (grey) interprets the interaction, reasons about the integration, and composes $\mathcal{P}_{\text{comp}}$ to guide the T2V generation.
+Inputs include: (1) a system prompt (red) guiding the VLM to produce VFX reasoning (purple), (2) a user prompt describing the content to integrate into the scene (green), and (3) key frames from the input video.
+The VLM produces VFX reasoning (purple) describing spatial placement, lighting, and motion for seamless integration, followed by $\mathcal{P}_{\text{comp}}$ (blue) captioning the target scene.
+Next, given the user prompt, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, the VLM identifies prominent elements from the original scene $\mathcal{O}_{\text{orig}}$ (yellow) and the new content to be added $\mathcal{O}_{\text{edit}}$ (brown).
+These outputs are used in our framework for the application of SAM, used to aid localization and harmonization of the new content.
+As can be seen in the output video, in this example, a majestic elephant is seamlessly integrated into the scene.
+Figure 17.
+System Prompt We Input the Pre-trained VLM (VFX Assistant).
+The system prompt contains guidelines to supply the VLM with the context of our task in order for it to interpret the user instruction, reason about the interaction, and accordingly, output a scene description that will be used for editing.
+Figure 18.
+System Prompt For Obtaining $\mathcal{O}_{\text{orig}}$ and $\mathcal{O}_{\text{edit}}$.
+Given the user prompt, the input video’s key frames and $\mathcal{P}_{\text{comp}}$, the VLM identifies prominent elements from the original scene $\mathcal{O}_{\text{orig}}$ and the new content to be added $\mathcal{O}_{\text{edit}}$.
+These outputs are used in our framework for the application of SAM, used to aid localization and harmonization of the new content.
+Figure 19.
+System Prompt for VLM-Based Evaluation.
+containing the guidelines for the evaluation setup, where original and edited frames are compared against an edit prompt across four criteria: prompt alignment, visual quality, content harmonization, and dynamics.